### v1

# Learning control variables and instruments for causal analysis in observational data

Abstract
This study introduces a data-driven, machine learning-based method to detect suitable control variables and instruments for assessing the causal effect of a treatment on an outcome in observational data, if they exist. Our approach tests the joint existence of instruments, which are associated with the treatment but not directly with the outcome (at least conditional on observables), and suitable control variables, conditional on which the treatment is exogenous, and learns the partition of instruments and control variables from the observed data. The detection of sets of instruments and control variables relies on the condition that proper instruments are conditionally independent of the outcome given the treatment and suitable control variables. We establish the consistency of our method for detecting control variables and instruments under certain regularity conditions, investigate the finite sample performance through a simulation study, and provide an empirical application to labor market data from the Job Corps study.
JEL Classification: C12, C21, C26
Keywords: treatment effects, causality, conditional independence, instrument, covariates, hypothesis test.
Addresses for correspondence: Nicolas Apfel, Murray Building (B58), Highfield Campus, University of Southampton, SO17 1TR Southampton, United Kingdom; n.apfel@soton.ac.uk. Julia Hatamyar, Research Fellow Centre for Health Economics, Heslington, York, YO10 5DD, United Kingdom; julia.hatamyar@york.ac.uk. Martin Huber, University of Fribourg, Bd. de Pérolles 90, 1700 Fribourg, Switzerland; martin.huber@unifr.ch. Jannis Kueck, Heinrich Heine University Düsseldorf, Universitätsstr. 1, 40225 Düsseldorf, Germany; kueck@dice.hhu.de.
1 Introduction
Methods for policy evaluation and causal analysis, aimed at quantifying the impact of a treatment or policy intervention on an outcome variable, conventionally rely on identifying assumptions considered untestable. For example, the well-known selection-on-observables, unconfoundedness, conditional independence, or ignorability assumption requires that the treatment is exogenous when conditioning on observed control variables, hereafter referred to as covariates, see for instance the survey by Imbens and Wooldridge (2009). The selection of covariates is typically justified based on theoretical and/or empirical reasoning, intuition, domain expertise, or prior empirical findings. Nonetheless, in many, if not most, empirical scenarios, such selections are subject to debate, given that the optimal set of covariates meeting the selection-on-observables assumption remains fundamentally uncertain.
In this paper, we suggest a machine learning-based procedure to simultaneously test the presence of (i) covariates satisfying the selection-on-observables assumption and (ii) relevant and valid instruments in observational data, as well as learning which variables in the data belong to either the set of covariates or instruments. When we refer to relevant and valid instruments, we mean variables that are associated with the treatment (relevance) but have no direct association with the outcome other than through the treatment (validity) conditional on covariates. We demonstrate that appropriate sets of covariates satisfying the identification requirements for treatment effects based on the selection-on-observables assumption, as well as relevant and valid instruments, can be detected in a data-driven way rather than being assumed by the researcher. For testing and learning covariates and instruments, we exploit a conditional independence condition that must hold when both relevant and valid instruments as well as appropriate covariates (satisfying the selection-on-observables assumption) exist: The instruments must be conditionally independent of the outcome, given the treatment and the covariates, see for instance the discussions in de Luna and Johansson (2014), Black et al. (2015), and Huber and Kueck (2022).
Our approach consists of the following steps. First, within the combined set of potential covariates and instruments, we sequentially test which variable is strongly associated with the treatment conditional on all remaining variables in that set. Second, we consider each of these strong predictors of the treatment as candidate instruments and sequentially test whether each of them is conditionally independent of the outcome when controlling for the treatment and all remaining variables in the combined set of potential covariates and instruments. If the conditional independence assumption is satisfied by (at least) one candidate instrument, then the instrument validity and selection-on-observables assumptions hold. This implies that the treatment is as good as random conditional on the remaining variables within the combined set of potential covariates and instruments. Consequently, treatment effects can be estimated utilizing methods that control for observed covariates, such as matching, regression, inverse probability weighting, or doubly robust techniques, as for instance discussed in Huber (2023).
Our discussion of the test focuses on conditional mean (rather than full) independence of the instrument, which implies the identification of average treatment effects (ATE). When assuming a limited set of observed variables (relative to the sample size), we employ regression for both selecting the candidate instruments in the first step and testing the conditional mean independence of the instrument in the second step. More concisely, testing is based on the mean squared difference in outcome prediction when regressing the outcome (1) on the treatment, the control variables, and the candidate instrument and (2) on the treatment and the control variables (but not the candidate instrument). This approach is closely related to the mean squared difference test based on a quadratic score function in Huber and Kueck (2022), but applies is sequentially to all candidate instruments. We demonstrate that our method is consistent for detecting covariates and instruments under certain regularity conditions. In a simulation study, we find support for this consistency result. However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations.
As an empirical illustration, we apply our method to data from an experimental study on Job Corps, a significant U.S. education program financed by the Department of Labor, targeting disadvantaged individuals aged 16 to 24, which has been previously analyzed in studies by Schochet et al. (2001) and Schochet et al. (2008). In this dataset, assignment to the program was randomized, making it a plausible candidate instrument for actual training participation in the first year of Job Corps, the treatment under investigation, if any direct effects of mere assignment on labor market outcomes like wage can be ruled out. Indeed, our approach selects random assignment as valid instrument. The method also suggests that training participation is exogenous conditional on a set of 28 pre-assignment covariates, with the p-value of the test statistic amounting to 76%.
Our paper contributes to a growing literature on testing identifying assumptions for causal inference. For instance, de Luna and Johansson (2014) and Black et al. (2015) make use of the same conditional independence of the instrument as considered here to test the selection-on-observables assumption for the treatment when assuming a valid instrument, based on matching or regression estimators. [^1] Angrist (2004), Brinch et al. (2017), and Huber (2013) consider related tests when instrument validity is holds unconditionally without controlling for covariates, in order to test the unconditional independence of the treatment and potential outcomes. Bertanha and Imbens (2015) consider the fuzzy regression discontinuity design (RDD), where instrument validity holds at a cutoff of a running variable which discontinuously affects treatment assignment. In contrast, Huber and Kueck (2022) jointly test both instrument validity and selection-on-observables assumptions with pre-defined instruments and covariates using machine learning approaches which permit for high dimensional covariates. This is conceptually closest to our approach, but one important difference is that the method of Huber and Kueck (2022) requires specifying the sets of supposed instruments and covariates to be used for testing, whereas in this paper, these sets of covariates and instruments are learned from the data and may therefore be a priori unknown.
Beyond that, we provide a new orthorgonalized quadratic score which improves on the quadratic score function used in Huber and Kueck (2022). This new score is of more general interest since it can be applied in many other contexts than ours as well. One example is the study by Parikh et al. (2024), who compare experimental and non-experimental treatment effect estimates and use an indicator for distinguishing between experimental and non-experimental evaluation designs, which has a comparable role as the instrument in our paper. Imposing the external validity of the experiment permits assessing the selection-on-observables assumption in the non-experimental design. Conversely, imposing the selection-on-observables assumption in the non-experimental design permits assessing the external validity of the experiment, which is equivalent to testing IV validity in our context. Unlike our paper, the authors do not consider testing both assumptions jointly. Also Angrist and Rokkanen (2015) use the same conditional independence of the instrument as considered here to test instrument validity within the framework of the sharp regression discontinuity design (RDD), where the selection-on-observables for the treatment holds by design, because it is a deterministic function of a cutoff in a running variable. This allows for testing whether the running variable is a valid instrument, i.e., whether it is not associated with the outcome conditional on the treatment. [^2] If the running variable is a valid instrument, causal effects can also be identified away from the cutoff, which is otherwise not feasible due to the lack of common support in the treatment across different values of the running variable. Again, our approach differs in that it tests the selection-on-observables and instrument validity assumptions jointly (rather than the latter conditional on the former).
Our study is also related to several contributions in the computer science-dominated literature on causal discovery, see see e.g. Kalisch and Bühlmann (2014), Peters et al. (2017), Glymour et al. (2019) for reviews. For instance, Peters et al. (2015) make use of pre-defined instruments and machine learning techniques to learn, in a data-driven way, which of the observed variables are treatments in the sense that they directly affect the outcome variable, under the assumption that these treatments satisfy the selection-on-observables assumption. The approach makes use of instruments in a way that may entail the rejection of treatments which violate the selection-on-observables assumptions, thereby providing power to detect identification failures. Our approach is different in that it focuses on a single, pre-defined treatment of interest, tests selection-on-observables and instrument validity jointly, and learns the sets covariates and instruments from the data.
Soleymani et al. (2022) and Quinzan et al. (2023) provide algorithms that select treatment variables in a data-driven way, while also controlling for observed covariates in a data-driven manner based on the double machine learning (DML) framework of Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey and Robins (2018). The idea is to sequentially consider each of the observed variables that may affect the outcome as treatment, while considering all remaining observed variables as covariates to estimate the direct effect of each candidate treatment on the outcome by DML. The algorithms retain only those observed variables as treatment variables that exhibit statistically significant effects on the outcome (where judging statistical significance should account for issues related to testing multiple hypotheses based on multiple candidate treatments). In contrast to our approach, these algorithms do not exploit instruments to test identifying assumptions, but assume the selection-on-observables assumption holds for all treatments. Nevertheless, the algorithms are useful also in our context, namely for selecting potential instruments with a sufficiently strong first stage effect on the pre-defined treatment conditional on other observed variables. [^3] Another domain of causal discovery related to our study is so-called Y-learning, see e.g. Mani et al. (2012) and Sevilla and Mayn (2021). Conditional on covariates, Y-learning implies that if two variables are independent of each other when not controlling for the treatment, statistically associated with each other when controlling for the treatment, and both independent of the outcome when controlling for the treatment, then these two variables are relevant and valid instruments. A further implication is that the selection-on-observables assumption holds. Our approach differs from Y-learning in that it imposes more causal structure by assuming that potential instruments and covariates are not affected by the treatment. For this reason, our method only requires a single (and a priori unknown) instrument, while Y-learning hinges on the existence of (at least) two instruments.
The computer science literature most closely related to our study is representation learning of sets of variables that are confounders jointly affecting a predefined treatment and outcome, instruments only affecting the treatment, and outcome predictors only affecting the outcome. Hassanpour and Greiner (2019) and Wu et al. (2021), for instance, propose deep learning algorithms minimizing global loss functions to simultaneously decompose pre-treatment variables into instruments, confounders, and outcome predictors. Yet, there are several differences between their studies and ours: First, their approaches impose selection-on-observables a priori to isolate instruments (whose exclusion from treatment effect estimation can increase efficiency) based on the same conditional independence condition as considered in our paper. However, we exploit the conditional independence condition to test selection-on-observables and instrument validity jointly. Second, their algorithms, which minimize a global loss function, aim at learning the full set of instruments, which is attractive for maximising efficiency in treatment effect estimation but might be very ambitious to achieve in a finite sample. In contrast, our algorithm of iteratively changing candidate instruments aims at a more moderate goal of detecting at least one instrument, which is a sufficient condition for identification. Third, we demonstrate the consistency of our algorithm in selecting valid instruments under certain regularity conditions, while the asymptotic behaviour of the deep learning methods in Hassanpour and Greiner (2019) and Wu et al. (2021) has not been derived.
Finally, our contribution also relates to a growing literature in statistics and econometrics employing machine learning techniques to separate valid from invalid instruments (which are confounders to be controlled for), see for instance Guo et al. (2018), Windmeijer et al. (2021), Windmeijer et al. (2019), Apfel and Liang (2021), and Apfel et al. (2022). These approaches rely on the assumption that a sufficiently large number of (a priori unknown) instruments is valid, and in order to detect the set of valid instruments Hansen-Sargan-type tests in combination with instrument-based estimators are used. In contrast, our approach does not pre-impose the existence of valid instruments, but tests the instrument validity (for a single instrument) and selection-on-observables assumptions jointly, in order to apply estimation based on the selection-on-observables assumption.
The remainder of this study is organized as follows. Section 2 discusses a set of identifying assumptions, including instrument relevance and validity as well as the selection-on-observables assumption, which imply the conditional independence of instruments and the outcome conditional on the treatment and covariates. Based on this conditional independence, Section 3 proposes machine learning-based procedures for jointly testing the instrumental variable and selection-on-observables assumptions in observational data. Section 4 suggests an algorithm that applies these tests sequentially when rotating the role of pre-treatment variables to detect relevant and valid instruments as well as covariate sets satisfying the selection-on-observables assumption in a data-driven manner. Section 5 provides a simulation study analyzing the finite sample performance of our method. Section 6 presents an application to labor market data from the Job Corps study. Section 7 concludes.
2 Identifying assumptions and testable conditional independence
This section reviews the findings in de Luna and Johansson (2014), Black et al. (2015), and Huber and Kueck (2022), which imply that under instrument validity and a selection-on-observables assumption concerning the treatment, the instrument is conditionally independent of the outcome given the treatment and observed covariates. To formalize the assumptions, let us denote by $D$ a treatment variable whose causal effect on an outcome variable $Y$ is of interest. Both $D$ and $Y$ might be discretely or continuously distributed. Using the potential outcomes framework as for instance advocated in Neyman (1923) and Rubin (1974), we denote by $Y(d)$ the potential outcome when exogenously setting the treatment $D$ of a subject to value $d$ in the support of the treatment. More generally, we will use capital and lower case letters for referring to random variables and specific values thereof, respectively. By representing the potential outcome $Y(d)$ as a function solely dependent on a subject’s own treatment status $D=d$, we implicitly adhere to the assumption that the potential outcomes of one subject are not influenced by the treatment status of others. This is known as the ‘stable unit treatment value assumption’ (SUTVA), see for instance the discussion in Rubin (1980) and Cox (1958), and is invoked throughout. Furthermore, we denote by $X$ and $Z$ sets of observed covariates and instruments, whose properties are yet to be defined. Based on this notation, we consider the same identifying assumptions as Huber and Kueck (2022).
The first assumption imposes some causal structure on our framework. First, it rules out the existence of reverse causality, meaning that the outcome cannot causally influence any other variables, and the treatment cannot causally affect any variables other than the outcome. We formalise this causal structure using the previously mentioned potential outcome notation, by applying the latter also to other variables. To this end, let $A(b)$ and $A(b,c)$ correspond to the potential value of variable $A$ when setting variable $B$ to $b$, or variables $B$ and $C$ to $b$ and $c$, respectively. Second, our first assumption enforces the principle of causal faithfulness, which dictates that variables linked by causal pathways, possibly conditional on other variables, are necessarily statistically dependent.
Assumption 1 (Causal structure).
$D(y)=D,\quad X(d,y)=X,\textrm{ and }Z(d,y)=Z\quad\forall d\in\mathcal{D}\textrm{ and }y\in\mathcal{Y},$
only variables which are d-separated in some causal model are statistically independent.
$\mathcal{D}$ and $\mathcal{Y}$ denote the support of $D$ and $Y$, respectively. The first line of Assumption 1 rules out a causal effect of outcome $Y$ on $D$, $X$, or $Z$ and of treatment $D$ on $X$ or $Z$. However, it allows for the possibility of both $X$ and $Z$ affecting $D$, $Y$, or even each other. This assumption aligns with the directed acyclic graph (DAG, see e.g. Pearl (2000)) presented in Figure 1, where causal relationships between variables are indicated by arrows: $Z$ and $X$ affect $D$, and $D$ and $X$ affect $Y$. Additionally, $X$ may influence $Z$ or vice versa, denoted by the bidirectional causal arrow. This is in line with the conventional practice of measuring covariates and instruments before treatment assignment, eliminating the potential for reverse causality between $D$ and $Y$ and the pre-treatment variables $X$ and $Z$. The DAG also features unobserved terms $U$ and $V$ that affect $Y$ and $D$, respectively, with dashed arrows denoting the unobservable nature of these effects. The second line of Assumption 1 enforces causal faithfulness, ensuring that only variables which are d-separated in the sense of Pearl (1988), i.e. not associated with each other via some causal paths (possibly conditional on other variables) are statistically independent (or conditionally independent).444 d-separation relies on blocking causal paths between variables. Formally, a path between two (sets of) variables $A$ and $B$ is blocked when conditioning on a (set of) control variable(s) $C$ if 1. the path between $A$ and $B$ is a causal chain, implying that $A\rightarrow M\rightarrow B$ or $A\leftarrow M\leftarrow B$, or a confounding association, implying that $A\leftarrow M\rightarrow B$, and variable (set) $M$ is among control variables $C$ (i.e. controlled for), 2. the path between $A$ and $B$ contains a collider, implying that $A\rightarrow S\leftarrow B$, and variable (set) $S$ or any variable (set) causally affected by $S$ is not among control variables $C$ (i.e. not controlled for). Based on this definition of blocking, the d-separation criterion states that $A$ and $B$ are d-separated when conditioning on control variable(s) $C$ if and only if $C$ blocks all paths between $D$ and $Y$. While d-separation is generally a sufficient condition for the (conditional) independence of two variables, it is a necessary condition under causal faithfulness.
Figure 1: Causal graph satisfying Assumption 1
The second assumption is a common support assumption concerning the treatment and the instrument:
Assumption 2 (Common support).
$\mathbb{P}(D=d,Z=z|X)>0\quad\forall d\in\mathcal{D}\textrm{ and }z\in\mathcal{Z},$
where $\mathcal{Z}$ denotes the support of $Z$. For discretely distributed treatments and instruments, Assumption 2 states that the joint probabilities of any $D=d$ and $Z=z$ conditional on $X$ are larger than zero. Under continuous treatment and/or instrument variables, then the joint probabilities are to be replaced by joint density functions conditional on $X$. Assumption 2 implies that both $\mathbb{P}(D=d|X)$, the so-called treatment propensity score, and $\mathbb{P}(Z=z|D,X)$, the instrument propensity score, are larger than zero. [^5] Assumption 2 is strictly speaking not required for testing, if testing only proceeds in a subgroup satisfying common support. However, this may imply reduced testing power.
The third assumption imposes a statistical association between the instrument and the treatment conditional on the covariates.
Assumption 3 (Conditional dependence between the treatment and instrument).
[…]
The fourth assumption invokes selection on observables, i.e. quasi-random treatment assignment conditional on covariates as e.g. considered in Imbens (2004):
Assumption 4 (Conditional independence of the treatment).
$Y(d){\perp\!\!\!\perp}D|X\quad\forall d\in\mathcal{D},$
[…]
The fifth assumption invokes instrument validity, requiring that the instrument is conditionally independent of the potential outcomes given the covariates $X$.
Assumption 5 (Conditional independence of the instrument).
$Y(d){\perp\!\!\!\perp}Z|X\quad\forall d\in\mathcal{D}.$
[…]
The previous assumptions can be used to construct tests for the identification of causal effects. Theorem 1 in Huber and Kueck (2022) demonstrates that conditional on Assumptions 1 and 3, the testable conditional independence $Y{\perp\!\!\!\perp}Z|D=d,X$ is necessary and sufficient for the joint satisfaction of Assumptions 4 and 5 when considering potential outcomes $Y(d)$ which match the factual treatment assignment $D=d$. Formally,
$Y(d){\perp\!\!\!\perp}D|X,\quad Y(d){\perp\!\!\!\perp}Z|X\iff Y{\perp\!\!\!\perp}Z|D=d,X\quad\forall d\in\mathcal{D}.$ (2.1)
Instead of verifying $Y{\perp\!\!\!\perp}Z|D,X$, we will focus on testing whether conditional mean independence of the instrument holds:
$\mathbb{E}[Y|D,X]=\mathbb{E}[Y|D,X,Z].$ (2.2)
Condition (2.2) is sufficient when considering the identification of average causal effects such as the conditional average treatment effect (CATE) given $X$, $\mathbb{E}[Y(1)-Y(0)|X]$, or the average treatment effect (ATE), $\mathbb{E}[Y(1)-Y(0)]$, rather than distributional parameters like quantile treatment effects. Theorem 2 in Huber and Kueck (2022) shows that (2.2) holds when replacing Assumptions 4 and 5 by the weaker conditional mean independence assumptions $\mathbb{E}[Y(d)|D,X]=\mathbb{E}[Y(d)|X]$ and $\mathbb{E}[Y(d)|Z,X]=\mathbb{E}[Y(d)|X]$ $\forall d\in\mathcal{D}$, as well as Assumption 3 by the first stage condition $\mathbb{E}[D|Z,X]\neq\mathbb{E}[D|X]$, implying that the conditional mean of $D$ varies with $Z$. Formally, conditional on Assumption 1 and $\mathbb{E}[D|X,Z]\neq\mathbb{E}[D|X]$, it holds that
$\mathbb{E}[Y(d)|D,X]=\mathbb{E}[Y(d)|X],\quad\mathbb{E}[Y(d)|X,Z]=\mathbb{E}[Y(d)|X]$
$\iff$ $\mathbb{E}[Y|D=d,X,Z]=\mathbb{E}[Y|D=d,X]\quad\forall d\in\mathcal{D}.$
The testable implication $\mathbb{E}[Y|D=d,X,Z]=\mathbb{E}[Y|D=d,X]$ is necessary and sufficient for the joint satisfaction of conditional mean independence of the treatment and the instrument when considering potential outcomes $Y(d)$ matching the (f)actual treatment assignment $D=d$.
Theorems 1 and 2 in Huber and Kueck (2022) imply that we are only able to test the respective selection-on-observables and instrument validity assumptions for factual outcomes, e.g. for the potential outcomes $Y(1)$ of individuals with $D=1$ and $Y(0)$ of individuals with $D=0$. Conversely, we cannot perform tests based on counterfactual outcomes, such as $Y(0)$ for individuals with $D=1$ and $Y(1)$ for individuals with $D=0$. Strictly speaking, we can for this reason only test a necessary, but not a sufficient condition for identification. Nevertheless, in practical terms, it seems unlikely that such violations would exclusively pertain to counterfactual outcomes, but never affect factual outcomes, because this would require highly specific modelling constraints. Therefore, testing the (mean) conditional independence of the instrument is likely to have the power to detect violations of the selection-on-observables and instrument validity assumptions in typical empirical applications.
The testing approach of Huber and Kueck (2022) requires the prior specification of the instrument $Z$ and covariates $X$. In contrast, our testing approach, as introduced below, does not require the predefinition of $Z$ and $X$ when testing (2.2). Instead, it learns them from the data by iteratively considering observed variables as instrument $Z$. This feature appears attractive in many practical contexts where obvious instruments are not available.
3 Testing based on double machine learning
We henceforth suggest a testing approach based on double machine learning based on the following null hypothesis $H_{0}$, which is equivalent to the conditional mean independence of the instrument $Z$ provided in condition (2.2):
$H_{0}:\mathbb{E}[Y|D=d,X=x,Z=z]-E[Y|D=d,X=x]=0\quad\forall d\in\mathcal{D},x\in\mathcal{X},z\ \in\mathcal{Z}.$ (3.1)
Under the null hypothesis $H_{0}$, the mean conditional outcome is constant across values of the instrument $Z$ given any value of $D$ and $X$, which may be tested for any values of $D$, $X$, and $Z$ in their respective support. However, if one or several variables are of rich support, this implies many testable implications - even infinitely many if one or more variables are continuous.
For this reason, one possible testing approach is to follow Huber and Kueck (2022) and test violations of (3.1) based on the mean squared difference between the conditional mean outcome when including versus excluding the instrument in the conditioning set. Denoting the conditional means by $\mu(d,x,z)=\mathbb{E}[Y|D=d,X=x,Z=z]$ and $m(D,X)=\mathbb{E}[Y|D=d,X=x]$, one aims at testing the following implication of equation (3.1):
$\mathbb{E}[\left(\mu(D,X,Z)-m(D,X)\right)^{2}]=0,$ (3.2)
based on a moment condition which uses the following Neyman (1959)-orthogonal score function:
$\phi(W,\theta,\eta)=(\eta_{1}(W)-\eta_{2}(W))^{2}-\theta+\zeta.$ (3.3)
$W=(Y,D,X,Z,\zeta)$ are random variables and $\eta=(\eta_{1},\eta_{2})$ are the so-called nuisance parameters, whose true values correspond to $\eta_{0,1}(W)=\mu(D,X,Z)$ and $\eta_{0,2}(W)=m(D,X)$. We note that the independent mean-zero random variable $\zeta$ in (3.3) is added to avoid a degenerate distribution of the estimator under $H_{0}$, a common problem in specification tests, see e.g. Hong and White (1995) and Wooldridge (1992). A disadvantage of testing based on the score function in equation (3.3) is the requirement to choose a random term $\zeta$, as the optimal selection of $\zeta$ in a given dataset is generally unknown. Further, while the estimator based on equation (3.3) is asymptotically normally distributed under the null hypothesis, as demonstrated in Huber and Kueck (2022), this is generally not the case under the alternative hypothesis. For this reason, we subsequently propose a new testing approach that is based on a refined score function that does not require user-selected random terms and entails a test statistic that is normally distributed under both the null and alternative hypotheses. [^6] We note that an alternative testing approach that could be applied in our context is suggested by Kook and Lundborg (2024)
To formally discuss the test, let us for the moment assume that the instrument $Z$ is binary and denote by $p(D,X)=\mathbb{P}(Z=1|D,X)$ the instrument propensity score. The score function considered in this case is given below:
$\quad\tilde{\psi}(W,\theta,\eta)$ (3.4)
$=(\mu(D,X,1)-\mu(D,X,0))^{2}$
$+2(\mu(D,X,1)-\mu(D,X,0))\left(\frac{(Y-\mu(D,X,1))\cdot Z}{p(D,X)}-\frac{(Y-\mu(D,X,0))\cdot(1-Z)}{1-p(D,X)}\right)$
$+\mu(D,X,1)-\mu(D,X,0)+\left(\frac{(Y-\mu(D,X,1))\cdot Z}{p(D,X)}-\frac{(Y-\mu(D,X,0))\cdot(1-Z)}{1-p(D,X)}\right)$
$-\theta,$
with $\theta_{0}=\mathbb{E}[\left(\mu(D,X,1)-\mu(D,X,0)\right)^{2}]+\mathbb{E}[\mu(D,X,1)-\mu(D,X,0)]$. Testing based on (3.4) corresponds to an aggregate $L_{2}$-type measure that can be used to test violations across values of $D$, $X$, and $Z$, which is common in specification tests based on nonparametric regression, see e.g. Racine (1997), Racine et al. (2006), Hong and White (1995) and Wooldridge (1992). In addition to the squared difference in conditional mean outcomes that we already encountered in the previous section, the score function notably contains a term in which the difference in conditional mean outcomes is multiplied with a difference in expressions obtained by inverse probability weighting (IPW) with the instrument propensity score. In fact, our new score above combines the orthogonalized squared score in (3.3) with the popular doubly robust score, as also considered for testing in Huber and Kueck (2022). Just as the squared difference in conditional mean outcomes, the score function $\tilde{\psi}$ is zero in expectations, $\mathbb{E}[\tilde{\psi}(W,\theta_{0},\eta_{0})]=0$, under the null hypothesis that $\theta_{0}=0$, which follows from iterated expectations, and satisfies the Neyman orthogonality property (see Appendix A.1).
In a next step, we adapt the score function in equation (3.4) to multivalued instruments $Z$, which is also a key contribution of this paper. In the case of a continuous $Z$, this requires discretizing its values in some parts of the score function. To this end, let $l=1,\dots,L$ be a partition of its support $\mathcal{Z}$ with $\cup_{l}Z_{l}=\mathcal{Z}$. For a discrete instrument, $Z_{l}=z_{l}$, $l=1,\dots,L$, would be any value $Z$ can take with probability $\mathbb{P}(Z=z_{l})>c$, with $c>0$. For a continuous instrument, such a partition may be generated based on the quantile function (e.g. percentiles) of $Z$. Let $1(Z\in Z_{l})$ denote the indicator function, which is one if $Z$ falls into the partition $Z_{l}$ and zero otherwise, and $p_{l}(D,X)=\mathbb{P}(Z\in Z_{l}|D,X)$ denote the corresponding instrument propensity score. Then, testing with a multivalued $Z$ can be based on the following score function:
$\quad\psi(W,\theta,\eta)$ (3.5)
$=\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))^{2}$
$+\sum_{l=1}^{L}2(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))$
$\left(\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)$
$+\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))$
$+\sum_{l=1}^{L}\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}-\theta.$
This score has a variance that is bounded away from zero, is zero in expectation under the null hypothesis and is Neyman-orthogonal, as formally shown in Appendix A.2. We may construct cross-fitted estimators of $\theta_{0}$ based on the score function (3.5). It is worth noting that the corresponding target parameter in (3.5) is given by
$\mathbb{E}\left[\sum_{l=1}^{L}(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))^{2}+(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right],$
which tests the null hypothesis given in equation (3.1) for binary and discrete instruments, and approximates equation (3.1) in the case of continuous instruments if the bins that define $Z_{l}$ become small.
When implementing the testing approach, we assume an i.i.d. sample of size $n$, in which $i$ denotes the index of an observation and $W_{i}=(Y_{i},D_{i},X_{i},Z_{i})$ are the variable values of observation $i$ in that sample, with $i$in $\{1,2,..,n\}$. When testing, we apply cross-fitting as for instance discussed in Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey and Robins (2018) to avoid over-fitting due to a correlation of the estimation of the nuisance parameters and $\theta_{0}$. Therefore, we split the data into $K$ subsamples of size $N=n/K$. The cross-fitted estimator is given by
$\hat{\theta}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{N,k}[\psi_{k}(W_{i},0,\hat{\eta})],$ (3.6)
with $\hat{\eta}=(\hat{\mu},\hat{m},\hat{p}_{1},\dots,\hat{p}_{L})$. Under the regularity conditions in Assumption 6, the estimator $\hat{\theta}$ is asymptotically normal and $\sqrt{n}$-consistent, as stated in Theorem 1. The proof is provided in Appendix A.3.
Assumption 6 (Asymptotic Normality).
The following assumption needs to hold for all $n\geq 3$, $\mathbb{P}\in\mathcal{P}$ and $q>2$: (i) $\|Y\|_{\mathbb{P},q}<C$ and $\mathbb{E}[\varepsilon^{2}1(Z\in Z_{l})]>c$ (ii) Given a random subset $I$ of $[n]$ of size $N=n/K$, the nuisance parameter estimator $\hat{\eta}_{0}=\hat{\eta}_{0}((W_{i})_{i\in I^{c}})$ obeys $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2q}\leq C$, $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},4}\leq\delta_{N}$, and $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2}\leq\delta_{N}^{1/2}N^{-1/4}$ with $\mathbb{P}$-probability not less than $1-o(1)$.
Theorem 1.
Conditional on Assumptions 6, the estimator in equation (3.6) satisfies
$\sqrt{n}\sigma^{-1}\hat{\theta}\rightsquigarrow N(\theta_{0},1),$ (3.7)
uniformly over $P\in\mathcal{P}$, where $\sigma^{2}=E[\psi(W,\theta_{0},\eta_{0})^{2}]$. Moreover, the result continues to hold if $\sigma^{2}$ is replaced by $\hat{\sigma}^{2}:=\mathbb{E}_{n}[(\psi(W_{i},\hat{\theta},\hat{\eta}))^{2}]$. Consequently, a test that rejects the null hypothesis $H_{0}$, $\theta_{0}=0$, if $|\sqrt{n}\hat{\sigma}^{-1}\hat{\theta}|>\Phi^{-1}(1-\alpha/2)$ has asymptotic level $\alpha$.
Theorem 1 states that the test statistic is normally distributed both under the null ($H_{0}:\theta_{0}=0$) and alternative hypothesis ($H_{1}:\theta_{0}\neq 0$). The following Corollary 1 shows that the proposed statistical test is consistent, i.e., the power converges to one if the sample size increases.
Corollary 1.
Let $c_{\alpha}:=\Phi^{-1}(1-\alpha/2)$ be the critical value of the test proposed above. Under the alternative hypothesis ($\theta_{0}\neq 0$), it holds
$\lim\limits_{n\rightarrow\infty}\,P(|\sqrt{n}\hat{\theta}/\hat{\sigma}|>c_{\alpha})=1.$
This holds true since
$P(|\sqrt{n}\hat{\theta}/\hat{\sigma}|>c_{\alpha})=P\left(\left|\widehat{\theta}_{j}\right|\geq c_{\alpha}\frac{\widehat{\sigma}_{j}}{\sqrt{n}}\right)\geq P\left(\left|\widehat{\theta}_{j}-\theta_{j}\right|\leq\left|\theta_{j}\right|-c_{\alpha}\frac{\widehat{\sigma}_{j}}{\sqrt{n}}\right)=1,$
as long as $c_{\alpha}=o(\sqrt{n})$.
4 Selection Method
The testing approaches outlined in the previous section 3 were conditional on having already defined the instrument $Z$ and covariates $X$ under which the selection-on-observables assumption with respect to the treatment supposedly holds. However, a main contribution of this study is a data-driven approach for learning partitions of observed pre-treatment variables into instruments and covariates. To do so, we suggest applying the testing approach iteratively when sequentially considering one variable from the set of all pre-treatment variables, henceforth denoted by $Q$, as instrument $Z$ and the remaining variables as covariates $X$. More specifically, our procedure consists of the following steps:
  1. 1.
    Select candidate instruments among observed variables $Q$ which are strongly correlated with $D$ conditional on the remaining variables. Let $\mathcal{S}$ denote the set of candidate instruments with first stage effects on treatment $D$ that are bounded away from zero conditional on the remaining variables in $Q$, which is to be estimated in the data based on some statistical criterion yet to be defined.
  2. 2.
    Iteratively define each variable in the estimated set of strong instruments, denoted by $\hat{S}$, as instrument $Z$ and all remaining variables in $Q$ as covariates $X$ when testing the hypothesis (3.2), e.g. based on the score functions in equations (3.4) or (3.5).
  3. 3.
    If testing suggests that the hypothesis (3.2) is not violated in one or several iterations according to a criterion of statistical significance yet to be defined, then select that candidate instrument in $\hat{S}$ for which the test statistic is closest to a non-violation as final instrument $Z$ and the remaining variables in $Q$ as final covariates $X$. If testing suggests that the hypothesis (3.2) is violated in all iterations, the testable implication (2.2) is rejected.
We note that the first step is required to select candidate instruments which, if ultimately selected as instrument $Z$, are statistically significant in terms of satisfying the first stage condition $E[D|X,Z]\neq E[D|X]$. Let us assume that $Q$ is low dimensional in the sense that the sample size $n$ is larger than the number of variables in $Q$, denoted by $p$, such that $n>p$. In this case, step one might be implemented based on a first stage regression of $D$ on $Q$ and selecting all regressors with statistically significant associations after controlling for multiple hypothesis testing issues into the set of candidate instruments $\hat{S}$.
As statistical criterion for selecting candidate instruments, we suggest first-stage hard thresholding, as performed in the IV selection literature for instance by Guo et al. (2018) and Masten and Poirier (2021). To formalize the discussion, let $F_{r}$ be the first stage F-statistic in the estimation of $E[D|X,Z]$ when defining $Z=Q_{r}$ and $X=Q_{[r]}$, with $Q_{[r]}=Q\setminus Q_{r}$ (and $Q=X\cup Z$). We obtain the estimated set of strong candidate instruments, denoted by $\hat{S}$, as those whose F-statistic exceeds a certain critical value $C_{n}$:
$\hat{\mathcal{S}}=\{r:F_{r}\geq C_{n}\}$ (4.1)
[…]
An alternative strategy is to iteratively consider each variable in $Q$ as instrument $Z$ and any remaining variables as covariates $X$ when estimating the first stage association $E[D|X,Z]$ based on methods for treatment or policy evaluation like doubly robust estimation, see e.g. Robins et al. (1994), Robins et al. (1995) and Chernozhukov et al. (2018). Such doubly robust approaches are also applicable in high-dimensional contexts with $n<<p$ when combining them with machine learning to control for important control variables among high-dimensional covariates $X$. For instance, the double machine learning algorithm suggested by Soleymani et al. (2022), selects direct causal parents in a data-driven manner by iteratively changing the definition of variables whose causal effect is of interest and the control variables. It may be applied to select the set of candidate instruments $S$ which are those variables in $Q$ strongly associated with $D$ conditional on the remaining elements in $Q$. Second, we proceed to test the null hypothesis in equation (3.2) iteratively over all candidate instruments that pass the first-stage threshold, $\hat{\mathcal{S}}$.
Our aim is to find a partition of variables for which the conditional independence of the respective candidate instrument holds. To discuss this more formally, we introduce the partition
$\mathcal{P}_{j}=\{Z=Q_{j},\ \ X=Q_{[j]}\},$
such that variable $j$ in set $Q$ is chosen to be the instrument, while the remaining variables in set $Q$ are used as control variables. Furthermore, let $\mathcal{V}$ denote the set of candidate instruments which are conditionally mean independent of the outcome, satisfying
$\mathcal{V}=\{j:E[Y|D,X]=E[Y|D,X,Z]\}.$
Furthermore, we denote the set of partitions for which the instrument has a first stage effect on the treatment and the conditional independence of the instrument holds by $\mathcal{P}^{*}$:
$\mathcal{P}^{*}=\{\mathcal{P}_{j}:j\in\left(\mathcal{S}\cap\mathcal{V}\right)\}.$ (4.2)
The corresponding estimated set of partition(s) $\mathcal{P}^{*}$, denoted by $\hat{\mathcal{P}}_{pass}$, is given by
$\hat{\mathcal{P}}_{pass}=\left\{\mathcal{P}_{j}:j\in\left(\hat{\mathcal{S}}\cap\hat{\mathcal{V}}\right)\right\},$ (4.3)
where $\hat{\mathcal{V}}=\{j:|\sqrt{n}\hat{\sigma}^{-1}_{j}\hat{\theta}_{j}|<c_{\alpha}\}$ is the set of instruments for which conditional independence is not rejected based on the test defined in Theorem 1 and Corollary 1 with significance level $\alpha$ and the corresponding critical value $c_{\alpha}$.
If $|\hat{\mathcal{P}}_{pass}|>1$ such that there is more than one single candidate instrument for which the first stage is sufficiently strong while conditional independence is not rejected, we can proceed in two ways. For a set of partitions $\mathcal{P}_{pass}=\{\mathcal{P}_{1},...,\mathcal{P}_{|\mathcal{P}_{pass}|},\}$ where each partition is associated with a specific IV, we define the set of IVs associated with that set of partitions as
=$\mathcal{I}=\left\{j:\mathcal{P}_{j}\in\mathcal{P}_{pass}\right\},$ (4.4)
we can then select the final partition into IVs and controls as
$\mathcal{P}_{all}=\{Z=Q_{\mathcal{I}},\ \ X=Q\setminus Q_{\mathcal{I}}\}.$ (4.5)
Alternatively, we can select the final partition as the one which maximizes the p-value when testing conditional independence:
$\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\hat{\mathcal{P}}_{pass}),$ (4.6)
where $p(\mathcal{P})$ is the $p$-value obtained in the conditional independence test of the instrument and the outcome. Even more crucial than the question how to choose the final partition if $|\hat{\mathcal{P}}_{pass}|>1$ is the question whether $\hat{\mathcal{P}}_{pass}=\emptyset$ or not.
The main contribution of this paper is the development of a new statistical procedure which tests the identification of a causal effect in a data-driven way. To this end, we consider the following hypotheses:
$H_{0}:\text{no identification}\quad v.s\quad H_{1}:\text{identification}\quad(Y(d){\perp\!\!\!\perp}D|X\quad\forall d\in\mathcal{D}).$ (4.7)
We conclude that $H_{1}$ is true, i.e, our causal effect of interest is identified if $\hat{\mathcal{P}}_{pass}\neq\emptyset$. The following Theorem 2 helps us to understand the type 1 error of our proposed test.
Theorem 2.
Under the assumptions of Theorem 1, assuming $\lim\limits_{n\rightarrow\infty}P(\hat{S}=\mathcal{S})=1$, for a given $\alpha$, it holds
$\lim\limits_{n\rightarrow\infty}P(\hat{\mathcal{P}}_{pass}\subseteq\mathcal{P}^{*})=1.$
Proof.
Since $\lim\limits_{n\rightarrow\infty}P(\hat{S}=\mathcal{S})=1$ we just have to show that
$\lim\limits_{n\rightarrow\infty}P(\hat{V}\subseteq\mathcal{V})=1.$
If $\hat{\mathcal{P}}_{pass}=\emptyset$, our statement holds trivially. Hence, consider a variable $j\in\hat{\mathcal{P}}_{pass}$, i.e, $Z_{j}\in\hat{\mathcal{S}}$ and $Z_{j}\in\hat{\mathcal{V}}$. Therefore, it holds
$|\sqrt{n}\hat{\sigma}^{-1}_{j}\hat{\theta}_{j}|<c_{\alpha}$
by definition of $\hat{\mathcal{V}}$. By Corollary 1, we have that
$\lim\limits_{n\rightarrow\infty}\,P(|\sqrt{n}\hat{\sigma}^{-1}_{j}\hat{\theta}_{j}|>c_{\alpha})=1$
if $j\notin\mathcal{V}$. This shows that $j\in\hat{\mathcal{V}}$ implies $j\in{\mathcal{V}}$ which concludes the proof. ∎
Theorem 2 states that $\lim\limits_{n\rightarrow\infty}P(\mathcal{V}\neq\emptyset)=1$ if $\hat{\mathcal{V}}\neq\emptyset$. Hence, the type 1 error of our proposed identification test in (4.7) goes to zero if sample size increases. This means that if our test finds identification ($\hat{\mathcal{P}}_{pass}\neq\emptyset$), there is identification with probability $1$ for large sample sizes. The next theorem provides us insights about the type 2 error of our test, i.e., how likely it is that we can find identification if there is identification. Theorem 3 states that the type 2 error is at least bounded by $\alpha$.
Theorem 3.
Assume that $\mathcal{P}^{*}\neq\emptyset$ (testable identification). Under the assumptions of Theorem 1, assuming $\lim\limits_{n\rightarrow\infty}P(\hat{S}=\mathcal{S})=1$, for a given $\alpha$, it holds
$\lim\limits_{n\rightarrow\infty}P(\hat{\mathcal{P}}_{pass}\neq\emptyset)\geq 1-\alpha.$
Proof.
Consider any variable $Z_{j}\in\mathcal{P}^{*}$. We know that $E[Y|D,X]=E[Y|D,X,Z_{j}]$ and hence by Theorem 1,
$P(\underbrace{|\sqrt{n}\hat{\sigma}_{j}^{-1}\hat{\theta}_{j}|>c_{\alpha}}_{:=A_{j}})=\alpha$
if $n\rightarrow\infty$. We can conclude that
$\lim\limits_{n\rightarrow\infty}P(\hat{\mathcal{P}}_{pass}\neq\emptyset)$ $=1-\lim\limits_{n\rightarrow\infty}P(\hat{\mathcal{P}}_{pass}=\emptyset)$
$=1-\lim\limits_{n\rightarrow\infty}P(\cap_{j=1,\dots,M}A_{j})\geq 1-\alpha$
with $M:=|\mathcal{P}^{*}|>0$. Understanding the stochastic dependence structure of the events $A_{j}$, $j=1,\dots,M$, could obviously lead to sharper bounds. ∎
Algorithm 1 describes the steps of our method by means of pseudo-code.
Algorithm 1 Testing procedure for selecting partition $\mathcal{P}^{*}$
1:procedure Testing
2:     Inputs: $Y\leftarrow Outcome$ $D\leftarrow Treatment$, $Q\leftarrow\textit{Potential controls and instruments}$
3:     Select strong instruments $\hat{\mathcal{S}}$:
4:     Regress: D $\leftarrow$ Q
5:     $\hat{F}_{j}\leftarrow\hat{t}_{j}^{2}$
6:     if $\hat{F}_{j}>C_{0.1/log(n)}$ then $j\in\hat{\mathcal{S}}$
7:     Select instruments passing conditional independence test $\hat{\mathcal{V}}$:
8:     for $j\in\hat{\mathcal{S}}$ do
9:         $\hat{\mu}(D,X,Z)\leftarrow$ and $\hat{m}(D,X)\leftarrow$ and $\hat{p}(D,X)\leftarrow$ estimates from machine learning
10:         $\hat{\theta}\leftarrow$ constructed via score in equation (3.4) (binary case) or (3.5) (continuous case)
11:         $\hat{t}_{ind,j}\leftarrow$ test $H_{0}:\hat{\theta}=0$
12:         if $\hat{t}_{ind,j}<t_{crit}$ then $j\in\hat{\mathcal{V}}$
13:     Select final partition $\hat{\mathcal{P}}$:
14:     $\hat{\mathcal{P}}_{pass}\leftarrow\{j\in\hat{\mathcal{S}},j\in\hat{\mathcal{V}}\}$
15:     if $\hat{\mathcal{P}}_{pass}=\emptyset$ then end and report that identification is “rejected”
16:     else if $|\hat{\mathcal{P}}_{pass}|=1$ then $\hat{\mathcal{P}}\leftarrow\hat{\mathcal{P}}_{pass}$
17:     else if $|\hat{\mathcal{P}}_{pass}|>1$ then
18:         $p(\hat{\mathcal{P}}_{pass})\leftarrow\text{p-value}(\hat{t}_{ind,j})$ for $j\in\hat{\mathcal{V}}$
19:         $\hat{\mathcal{P}}=\underset{\mathcal{P}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P})$
20:     In case identification is confirmed ($\hat{\mathcal{P}}_{pass}\neq\emptyset$), estimate the treatment effect:
21:     Regress: $Y\leftarrow D$ and $X=Q_{[j]}\in\hat{\mathcal{P}}$
5 Simulation study
This section provides a simulation study to investigate the finite sample behavior of our testing approach based on the following data generating process:
$Y$ $=$ $D+X^{\prime}\beta+\gamma Z+W+U,$
$D$ $=$ $I\{X^{\prime}\beta+Z+\delta W+V>0\},$
$X,Z$ $\sim$ $binomial(\pi),$
$W$ $\sim$ $N(0,1),U\sim N(0,1),V\sim N(0,1),$
with $X,Z,W,U,V$ being independent of each other. Outcome $Y$ is a linear function of $D$ (whose treatment effect is one), covariates $X$ (for $\beta\neq 0$), the unobservables $W$ and $U$, and the supposed instrument $Z$ if the coefficient $\gamma\neq 0$. The binary treatment $D$ is a function of $X$ and the unobservable $V$, as well as $W$ if coefficient $\delta\neq 0$. While the supposed instrument $Z$ is binary, the unobserved terms $U,V,W$ are random, standard normally distributed variables that are independent of each other, of $Z$, and of $X$. Covariates $X$ are created as follows. We first draw a matrix of multivariate normal variables $\tilde{X}\sim N(\mathbf{0},\bm{\Sigma})$ where $\bm{\Sigma}$ is a covariance matrix with $Cov(X_{j},X_{k})=0.5^{|j-k|}$. We then compute probabilities $\pi_{i}=\frac{1}{1+exp(-\tilde{X}_{i})}$ and draw $X_{i}\sim Bernoulli(\pi_{i})$. $\beta$ gauges the effects of the covariates on $Y$ and $D$, respectively, and thus, the magnitude of confounding due to observables. The $j$th element in the coefficient vector $\beta$ is set to $0.8/j$ for $j=1,...,4$, implying a linear decay of covariate importance in terms of confounding for the first 4 covariates. We use the final element, $j=20$ as the instrument $Z$. This results in 20 total covariates, with the first 4 as confounders, and the last as the instrument $Z$.
We investigate the performance of our testing approach in $100$ simulations with sample sizes of $n=1000$, $4000$, and $16000$, setting the number of covariates to $20$. Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with two folds ($K=2$) when using the doubly robust score function in equation (3.4). For nuisance parameter estimation, lasso regression with default parameters as implemented in the glmnet package by Friedman et al. (2010) for the statistical software R is applied. We consider several statistics from our simulations: the estimated violation, $\hat{\theta}$, when using variable $Z$ as the instrument (est), the standard deviation of these estimates (std), and the average of the standard errors across all samples (mean se). These statistics are useful for judging the performance of the test conditional on selecting $Z$ as the instrument, which is the correct choice if $\gamma=0$, while there is no valid instrument if $\gamma\neq 0$.
However, since the choice of the instrument is not fixed a priori but is part of the estimation process, we also consider the following two performance measures. The first one is the empirical detection rate of $Z$ (det.Z), which indicates the frequency with which variable $Z$ is selected as the best candidate instrument and has a p-value greater than 10% when testing conditional independence based on the estimate $\hat{\theta}$. In other words, this corresponds to the proportion of simulations in which our method simultaneously selects $Z$ as the instrument and keeps the null hypothesis. In scenarios where the null hypothesis holds ($\delta=\gamma=0$), this detection rate should approach one as the sample size increases. The second measure is an equivalent detection rate for the confounders $X_{1}$ to $X_{4}$ (det.X). It indicates the frequency with which one of the confounders is selected as the best candidate instrument and yields a p-value greater than 10% when testing conditional independence using this confounder as the instrument. As the sample size increases, this detection rate should always approach zero, because confounders are never valid instruments.
Table 1: Simulations: Binary Instrument
N est std mean se det.Z det.X
Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
1000 0.002 0.039 0.018 31% 48%
4000 -0.001 0.010 0.006 46% 27%
16000 -0.000 0.003 0.002 61% 0%
Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
1000 -0.027 0.029 0.019 39% 21%
4000 -0.027 0.014 0.010 8% 21%
16000 -0.026 0.006 0.006 0% 0%
Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$)
1000 -0.353 0.185 0.146 6% 69%
4000 -0.457 0.111 0.091 0% 72%
16000 -0.578 0.072 0.053 0% 2%
Notes: columns ‘est’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z$ as instrument, its standard deviation, and the average of the standard errors across all samples, respectively. ‘det.Z’ gives the empirical detection rate of selecting $Z$ as instrument and obtaining a p-value $>0.10$ (or 10%) when estimating $\theta$. This corresponds to the frequency with which $Z$ is selected as the best candidate instrument and also entails a p-value of more than 10% when testing conditional independence. ‘det.X’ is an equivalent measure for confounders $X_{1}$ to $X_{4}$. It corresponds to the frequency with which one of the confounders is selected as the best candidate instrument and also entails a p-value of more than 10% when testing conditional independence.
Table 1 reports the simulation results for a binary $Z$ and binary $X_{j}$. $Z$ is correlated with $X_{j}$ according to the covariance matrix $\bm{\Sigma}$. The top rows focus on the case where $\delta=\gamma=0$, meaning that both Assumptions 4 and 5 are satisfied. As the sample size increases, the test is more likely to simultaneously select the true instrument $Z$ and keep the null hypothesis that the assumptions are satisfied. Specifically, the detection rate of $Z$ (det.Z) increases from 31% with the smaller sample size of 1000 observations to 61% with 16000 observations. This improvement is mirrored by a reduction in the rate at which a confounder variable ($X_{1}...X_{4}$) is selected as the instrument and passes the test (det.X). Under the largest sample size of 16000 observations, it never occurs that any confounder is selected as the best instrument and passes the conditional independence test. Additionally, we observe that, conditional on having selected the true instrument $Z$, the estimated violation $\hat{\theta}$ (est) is close to zero for any sample size.
The intermediate rows present the results for a violation of Assumption 4, selection on observables, when setting $\delta=2$, $\gamma=0$. As the sample size increases, the detection rates of $Z$ (det.Z) and the confounder variables (det.X) both go to zero, as expected. Furthermore, the estimated violation $\hat{\theta}$ conditional on using $Z$ as an instrument (est) is now bounded away from zero across all samples. The lower rows of Table 1 provide the performance measures under a violation of Assumption 5, instrument validity, when considering $\delta=0$, $\gamma=0.5$. Again and as it should be the case, both the detection rates of $Z$ (det.Z) and the confounder variables (det.X) tend to zero as the sample size increases. At the same time, the estimated violation conditional on using $Z$ as an instrument (est) is substantially biased across all samples.
Our simulations point to the consistency of our procedure in terms of selecting the correct instrument when Assumptions 4 and 5 are satisfied, and the rejection of the null hypothesis when violations of these assumptions occur. However and as a word of caution for empirical applications, our simulations also show that the sample size might need to be quite substantial for our testing approach to perform adequately, as the finite sample performance in our simulations is not convincing in moderate samples of just several thousand observations.
6 Empirical application
This section provides an empirical application to the Job Corps (JC) training program, one of the largest and most comprehensive job training programs for disadvantaged youth in the US that offers various types of academic and vocational training. The data come from the National Job Corps Study, a randomized experiment conducted in the mid-to-late 1990s in the United States to evaluate the effectiveness of JC on various labour market outcomes. Schochet et al. (2001) and Schochet et al. (2008) evaluate the impact of random program assignment on a wide range of labor market outcomes and find positive effects on education, employment, and earnings in the longer run.
We employ our testing methodology to analyze the experimental JC data as available in the causalweight package for the statistical software R, provided by Bodory and Huber (2018). This dataset named JC comprises 9240 observations and 46 variables. Among these variables is a binary indicator denoting participation in any form of academic or vocational training within the first year following program assignment, which serves as our treatment variable of interest $D$. It is noteworthy that while assignment to JC is randomized, actual participation in training might be selective due to non-compliance with the assignment. For instance, some individuals eligible for JC may choose not to participate. Our outcome variable are weekly earnings in the fourth year post-JC assignment. Alongside random program assignment, which stands as an obvious candidate instrument, our vector $Q$ encompasses 28 pre-assignment characteristics. These include gender, age, ethnicity, education level, mother tongue, marital status, a binary indicator for having at least one child, previous labor market participation and earnings, household size, parents’ education, welfare receipt during childhood, as well as variables related to health and health behavior. Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with two folds ($K=2$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively. In the case of a continuous candidate instrument, its support is partitioned based on the quartiles of its distribution, such that $L=4$. As in the simulations, we apply lasso regression for nuisance parameter estimation.
When running our algorithm, it selects random program assignment as the instrument, yielding a very high p-value of 76% when running the conditional independence test. On the one hand, this indicates the validity of the instrumental variable (IV), implying not only the randomness of program assignment inherent to the experimental design. It also suggests that the assignment does not directly affect the earnings outcome other than through the treatment (e.g. through motivation or disappointment when being or not being eligible for the program). On the other hand, the testing result suggests that the selection-on-observables assumption holds for the treatment when controlling for the pre-assignment covariates available in the data. Consequently, given the covariates, our result suggests that we may evaluate the average treatment effect (ATE) on the total population. In contrast, employing an IV-based approach utilizing random assignment as the instrument for effect estimation would, under specific additional assumptions like treatment monotonicity in the instrument, only permit assessing the local average treatment effect (LATE) on the subpopulation of compliers, whose participation in training aligns with the random assignment, as discussed in Imbens and Angrist (1994) and Angrist et al. (1996).
7 Conclusion
In this paper, we introduced a machine learning-based algorithm based on a novel doubly robust score function designed to detect and test, in a data-adaptive manner, the presence of control variables sufficient for identifying treatment effects in observational data, as well as variables satisfying instrument validity. The method learns the partition of observed variables into instruments and control variables from the data, relying on the condition that valid instruments are conditionally independent of the outcome given the treatment and control variables that imply the conditional treatment exogeneity. This represents an important advancement over previously suggested tests of identifying assumptions for causal inference in observational data, which rely on a priori assumptions about whether a variable is a control variable or an instrument.
We demonstrated the consistency of the method for detecting control variables and instruments (if they exist) under certain regularity conditions and investigated its finite sample performance through a simulation study, which supports the consistency result. However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations. Finally, we applied our algorithm to empirical labor market data from the Job Corps study. Our approach (as expected) selected random assignment into the Job Corps program as valid instrument, while also suggesting that training participation in the first year after assignment, our treatment of interest, is exogenous conditional on pre-assignment covariates.
Appendix A Appendix
A.1 Proof of moment condition and Neyman orthogonality of $\tilde{\psi}$
Equation (3.4) suggests the following score function for testing when $Z$ is binary:
$\quad\tilde{\psi}(W,\theta,\eta)$ (A.1)
$=(\mu(D,X,1)-\mu(D,X,0))^{2}$
$+2(\mu(D,X,1)-\mu(D,X,0))\left(\frac{(Y-\mu(D,X,1))\cdot Z}{p(D,X)}-\frac{(Y-\mu(D,X,0))\cdot(1-Z)}{1-p(D,X)}\right)$
$+\mu(D,X,1)-\mu(D,X,0)+\left(\frac{(Y-\mu(D,X,1))\cdot Z}{p(D,X)}-\frac{(Y-\mu(D,X,0))\cdot(1-Z)}{1-p(D,X)}\right)$
$-\theta$ (A.2)
$:=\tilde{\psi}_{1}(W,\theta,\eta)+\tilde{\psi}_{2}(W,\theta,\eta)-\theta$
with
$\quad\tilde{\psi}_{1}(W,\theta,\eta)$
$=(\mu(D,X,1)-\mu(D,X,0))^{2}$
$+2(\mu(D,X,1)-\mu(D,X,0))\left(\frac{(Y-\mu(D,X,1))\cdot Z}{p(D,X)}-\frac{(Y-\mu(D,X,0))\cdot(1-Z)}{1-p(D,X)}\right)$
and
$\tilde{\psi}_{2}(W,\theta,\eta)=\mu(D,X,1)-\mu(D,X,0)+\left(\frac{(Y-\mu(D,X,1))\cdot Z}{p(D,X)}-\frac{(Y-\mu(D,X,0))\cdot(1-Z)}{1-p(D,X)}\right).$
The moment condition $\mathbb{E}[\tilde{\psi}(W,\theta_{0},\eta_{0})]=0$ holds, because
$\mathbb{E}\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))^{2}+(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\right]-\theta=0$
and
$\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\right]=0,$
as
$\mathbb{E}\left[\frac{(Y-\mu_{0}(D,X,1))Z}{p_{0}(D,X)}\right]$ $=\mathbb{E}\left[\frac{Z}{p_{0}(D,X)}\mathbb{E}\left[Y-\mu_{0}(D,X,1)|D,X,Z\right]\right]$
$=P(Z=1)\mathbb{E}\left[\frac{1}{p_{0}(D,X)}\left(\mathbb{E}\left[Y|D,X,Z\right]-\mu_{0}(D,X,1)\right)\bigg{|}Z=1\right]$
$=P(Z=1)\mathbb{E}\left[\frac{1}{p_{0}(D,X)}\left(\mathbb{E}\left[Y|D,X,Z=1\right]-\mu_{0}(D,X,1)\right)\bigg{|}Z=1\right]=0$
and analogously $\mathbb{E}\left[\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right]=0$. Furthermore, by the same argument, we have
$\mathbb{E}\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\right]=0.$
Neyman orthogonality of $\tilde{\psi}$ can be shown by taking the Gateaux derivates w.r.t. the nuisance parameters:
$\quad\partial_{r}E[\tilde{\psi}_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
$=E\left[\partial_{r}\tilde{\psi}_{1},(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big{|}_{r=0}\right]$
$=2E\left[((\mu_{0}(D,X,1)-\mu_{0}(D,X,0))((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))\}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(Y-\mu_{0}(D,X,1))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
$\quad+2E\bigg{[}((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))$
$\quad\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\bigg{]}$
$=2E\left[((\mu_{0}(D,X,1)-\mu_{0}(D,X,0))((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))\}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$=2E\left[((\mu_{0}(D,X,1)-\mu_{0}(D,X,0))((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))\}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,1)-\mu_{0}(D,X,1))\mathbb{E}\left[\frac{Z}{p_{0}(D,X)}\Bigg{|}D,X\right]\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,0)-\mu_{0}(D,X,0))\mathbb{E}\left[\frac{(1-Z)}{1-p_{0}(D,X)}\Bigg{|}D,X\right]\right]$
$\quad=0$
and
$\quad\partial_{r}\mathbb{E}[\tilde{\psi}_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
$=\mathbb{E}\left[\partial_{r}\tilde{\psi}_{2},(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big{|}_{r=0}\right]$
$=\mathbb{E}\left[(\mu(D,X,1)-\mu_{0}(D,X,1))\right]-\mathbb{E}\left[(\mu(D,X,0)-\mu_{0}(D,X,0))\right]$
$\quad-\mathbb{E}\left[\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\quad+\mathbb{E}\left[\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$\quad-\mathbb{E}\left[\frac{Z(Y-\mu_{0}(1,D,X))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
$\quad-\mathbb{E}\left[\frac{(1-Z)(Y-\mu_{0}(0,D,X))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]=0$
since $\mathbb{E}[Z|D,X]=p_{0}(D,X)$, $\mathbb{E}[Z(Y-\mu_{0}(D,X,1))|D,X]=0$ and $\mathbb{E}[(1-Z)(Y-\mu_{0}(D,X,0))|D,X]=0$.
A.2 Proof of moment condition and Neyman orthogonality of $\psi$
Equation (3.5) suggests the following score function for testing when $Z$ is multivalued discrete or continuous:
$\quad\psi(W,\theta,\eta)$ (A.3)
$=\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))^{2}$
$+\sum_{l=1}^{L}2(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))$
$\left(\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)$
$+\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))$
$+\sum_{l=1}^{L}\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}-\theta$
$:=\psi_{1}(W,\theta,\eta)+\psi_{2}(W,\theta,\eta)-\theta.$
First, we show that the moment condition holds. By definition, we have
$\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))^{2}+\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))-\theta=0.$
Analogous to the proof in Appendix A.1, we have
$\sum_{l=1}^{L}\mathbb{E}\left[\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right]=0.$
Hence, we have to show that
$\mathbb{E}\Bigg{[}\sum_{l=1}^{L}(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))$
$\quad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\Bigg{]}=0,$
which holds by the same argument. Next, we show that Neyman orthogonality holds. First, note that
$\quad\partial_{r}\mathbb{E}[\psi_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
$=\mathbb{E}\left[\partial_{r}\psi_{2},(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big{|}_{r=0}\right]$
$=\sum_{l=1}^{L}\Bigg{(}\mathbb{E}\left[(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]-\mathbb{E}\left[(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
$\quad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$\quad+\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$\quad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$\quad-\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]\Bigg{)}=0$
since
$\quad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$=\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]=0,$
by the same arguments used before,
$\quad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$=\sum_{l=1}^{L}\mathbb{E}\left[\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\Bigg{|}D,X\right]\right]$
$=\sum_{l=1}^{L}\mathbb{E}\left[(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\mathbb{E}\left[\frac{1(Z\notin Z_{l})}{1-p_{l}(D,X)}\Bigg{|}D,X\right]\right]$
$=\sum_{l=1}^{L}\mathbb{E}\left[(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
by iterated expectation and
$\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$=\sum_{l=1}^{L}\mathbb{E}\left[(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right].$
$\quad\partial_{r}\mathbb{E}[\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))]\big{|}_{r=0}$
$=\mathbb{E}\left[\partial_{r}\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big{|}_{r=0}\right]$
$=2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\bigg{[}((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l})))$
$\quad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))\cdot 1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))\cdot 1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\bigg{]}$
$=2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
$=0.$
A.3 Proof of Theorem 1
To prove Theorem 1, we apply Theorem 3.1 in Chernozhukov et al. (2018). Hence, we only need to show Assumption 3.1 and 3.2 in Chernozhukov et al. (2018). All bounds in the proof hold uniformly over $\mathbb{P}\in\mathcal{P}$ but we omit this qualifier for brevity. We use $C$ to denote a strictly positive constant that is independent of $n$ and $\mathbb{P}\in\mathcal{P}$. The value of $C$ may change at each appearance. In Appendix A.2 we have already shown that the moment condition and Neyman orthogonality is satisfied. Next, we note that the score in equation (3.5) is linear, i.e.
$\psi(W,\theta,\eta)=\psi^{a}(W,\eta)\theta+\psi^{b}(W,\eta),$
with $\psi^{a}(W,\eta)=-1,$ which is in line with Assumption 3.1 in Chernozhukov et al. (2018). Next, we demonstrate the satisfaction of Assumption 3.2 to complete the proof. We define the following nuisance realization set $\mathcal{T}_{n}$ as the set of all P-square-integrable functions $\eta$ such that
$\|\eta_{0}-\eta\|_{P,2q}$ $\leq C,$
$\|\eta_{0}-\eta\|_{P,4}$ $\leq\delta_{N},$
$\|\eta_{0}-\eta\|_{P,2}$ $\leq\delta_{N}^{1/2}N^{-1/4},$
for $\delta_{N}=o(1)$ and a constant $q>2$. Note that Assumption 3.2 (a)-(c) hold by construction of the set $\mathcal{T}_{N}$ and Assumption 6 due to same arguments as in Huber and Kueck (2022). Assumption 3.2 (d) in Chernozhukov et al. (2018) holds since
$\quad\mathbb{E}[\psi(W,\theta_{0},\eta_{0})^{2}]$
$\geq\mathbb{E}\left[\left(\sum_{l=1}^{L}\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)^{2}\right]$
$>0$
by Assumption 6 and since $p_{l}(D,X)>c>0$ by construction for all $l=1,\dots,L$. This completes the proof.

### v2

@@
 Abstract
-This study introduces a data-driven, machine learning-based method to detect suitable control variables and instruments for assessing the causal effect of a treatment on an outcome in observational data, if they exist.
+This study introduces a data-driven, machine learning-based method to detect suitable control variables and instruments for assessing the causal effect of a treatment on an outcome in observational data.
 Our approach tests the joint existence of instruments, which are associated with the treatment but not directly with the outcome (at least conditional on observables), and suitable control variables, conditional on which the treatment is exogenous, and learns the partition of instruments and control variables from the observed data.
 The detection of sets of instruments and control variables relies on the condition that proper instruments are conditionally independent of the outcome given the treatment and suitable control variables.
-We establish the consistency of our method for detecting control variables and instruments under certain regularity conditions, investigate the finite sample performance through a simulation study, and provide an empirical application to labor market data from the Job Corps study.
+We establish the consistency of our method for detecting control variables and instruments under certain regularity conditions, investigate the finite sample performance through a simulation study, and provide an empirical application to health data from the Oregon Health Insurance Experiment.
 JEL Classification: C12, C21, C26
 Keywords: treatment effects, causality, conditional independence, instrument, covariates, hypothesis test.
-Addresses for correspondence: Nicolas Apfel, Murray Building (B58), Highfield Campus, University of Southampton, SO17 1TR Southampton, United Kingdom; n.apfel@soton.ac.uk.
+We have benefited from comments by Niels Richard Hansen and Leonard Henckel.
+Addresses for correspondence: Nicolas Apfel, Department of Economics Faculty of Economics and Statistics, University of Innsbruck, Universitaetsstrasse 15, 6020 Innsbruck, nicolas.apfel@uibk.ac.at.
 Julia Hatamyar, Research Fellow Centre for Health Economics, Heslington, York, YO10 5DD, United Kingdom; julia.hatamyar@york.ac.uk.
@@
 More concisely, testing is based on the mean squared difference in outcome prediction when regressing the outcome (1) on the treatment, the control variables, and the candidate instrument and (2) on the treatment and the control variables (but not the candidate instrument).
-This approach is closely related to the mean squared difference test based on a quadratic score function in Huber and Kueck (2022), but applies is sequentially to all candidate instruments.
+This approach builds on the mean squared difference test based on a quadratic score function in Huber and Kueck (2022), but applies it sequentially across all candidate instruments.
+Unlike the original test, our procedure reverses the hypothesis structure: the null hypothesis states that identification fails, while the alternative hypothesis indicates successful identification.
 We demonstrate that our method is consistent for detecting covariates and instruments under certain regularity conditions.
-In a simulation study, we find support for this consistency result.
-However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations.
-As an empirical illustration, we apply our method to data from an experimental study on Job Corps, a significant U.S.
-education program financed by the Department of Labor, targeting disadvantaged individuals aged 16 to 24, which has been previously analyzed in studies by Schochet et al.
-(2001) and Schochet et al.
+In a simulation study with ten covariates, we find support for this consistency result.
+However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset.
+As an empirical illustration, we apply our method to health data from the Oregon Health Insurance Experiment, previously analysed by Finkelstein et al.
 (2012).
-In this dataset, assignment to the program was randomized, making it a plausible candidate instrument for actual training participation in the first year of Job Corps, the treatment under investigation, if any direct effects of mere assignment on labor market outcomes like wage can be ruled out.
-Indeed, our approach selects random assignment as valid instrument.
-The method also suggests that training participation is exogenous conditional on a set of 28 pre-assignment covariates, with the p-value of the test statistic amounting to 76%.
+In this experiment, low-income adults were randomly assigned the opportunity to apply for Medicaid, a public health insurance program in the United States.
+The random assignment provides a plausible instrumental variable for actual Medicaid enrolment - the treatment of interest - provided that assignment itself has no direct effect on health outcomes such as doctor visits.
+Our approach indeed selects random assignment as a valid instrument, together with several other variables, and indicates that Medicaid enrolment is exogenous, conditional on a rich set of more than 200 pre-assignment covariates capturing a wide range of socio-economic and health-related characteristics.
 Our paper contributes to a growing literature on testing identifying assumptions for causal inference.
@@
 (2015) make use of the same conditional independence of the instrument as considered here to test the selection-on-observables assumption for the treatment when assuming a valid instrument, based on matching or regression estimators.
-[^1] Angrist (2004), Brinch et al.
-(2017), and Huber (2013) consider related tests when instrument validity is holds unconditionally without controlling for covariates, in order to test the unconditional independence of the treatment and potential outcomes.
+[^1] See also the related test by Chen et al.
+(2018), which (in contrast to other methods) requires symmetrically distributed error terms.
+Furthermore, Angrist (2004), Brinch et al.
+(2017), and Huber (2013) assume instrument validity to hold unconditionally without controlling for covariates, in order to test the unconditional independence of the treatment and potential outcomes.
 Bertanha and Imbens (2015) consider the fuzzy regression discontinuity design (RDD), where instrument validity holds at a cutoff of a running variable which discontinuously affects treatment assignment.
-In contrast, Huber and Kueck (2022) jointly test both instrument validity and selection-on-observables assumptions with pre-defined instruments and covariates using machine learning approaches which permit for high dimensional covariates.
+In contrast, Huber and Kueck (2022) jointly test both instrument validity and selection-on-observables assumptions with pre-defined instruments and covariates using machine learning approaches which permit for high-dimensional covariates.
+[^2] A joint test is also proposed by Angrist et al.
+(2017), however, for parametric regression models with low dimensional covariates.
 This is conceptually closest to our approach, but one important difference is that the method of Huber and Kueck (2022) requires specifying the sets of supposed instruments and covariates to be used for testing, whereas in this paper, these sets of covariates and instruments are learned from the data and may therefore be a priori unknown.
-Beyond that, we provide a new orthorgonalized quadratic score which improves on the quadratic score function used in Huber and Kueck (2022).
-This new score is of more general interest since it can be applied in many other contexts than ours as well.
+Beyond that, we provide a new orthogonalized quadratic score, which improves upon the quadratic score function used in Huber and Kueck (2022).
+This new score is of broader interest since it can be applied in many other contexts beyond ours.
 One example is the study by Parikh et al.
-(2024), who compare experimental and non-experimental treatment effect estimates and use an indicator for distinguishing between experimental and non-experimental evaluation designs, which has a comparable role as the instrument in our paper.
-Imposing the external validity of the experiment permits assessing the selection-on-observables assumption in the non-experimental design.
-Conversely, imposing the selection-on-observables assumption in the non-experimental design permits assessing the external validity of the experiment, which is equivalent to testing IV validity in our context.
+(2024), who compare experimental and non-experimental treatment effect estimates and use an indicator to distinguish between experimental and non-experimental evaluation designs, which plays a role comparable to the instrument in our paper.
+Imposing the external validity of the experiment allows for assessing the selection-on-observables assumption in the non-experimental design.
+Conversely, imposing the selection-on-observables assumption in the non-experimental design allows for assessing the external validity of the experiment, which is equivalent to testing IV validity in our context.
 Unlike our paper, the authors do not consider testing both assumptions jointly.
-Also Angrist and Rokkanen (2015) use the same conditional independence of the instrument as considered here to test instrument validity within the framework of the sharp regression discontinuity design (RDD), where the selection-on-observables for the treatment holds by design, because it is a deterministic function of a cutoff in a running variable.
-This allows for testing whether the running variable is a valid instrument, i.e., whether it is not associated with the outcome conditional on the treatment.
+Similarly, Angrist and Rokkanen (2015) employ the same conditional independence of the instrument as considered here to test instrument validity within the framework of the sharp regression discontinuity design (RDD), where selection-on-observables for the treatment holds by design, as it is a deterministic function of a cutoff in a running variable.
+This allows testing whether the running variable is a valid instrument, i.e., whether it is not associated with the outcome conditional on the treatment.
 [^3] If the running variable is a valid instrument, causal effects can also be identified away from the cutoff, which is otherwise not feasible due to the lack of common support in the treatment across different values of the running variable.
-Again, our approach differs in that it tests the selection-on-observables and instrument validity assumptions jointly (rather than the latter conditional on the former).
+Relatedly, one strand of the statistics literature imposes the selection-on-observables assumption in observational studies and exploits the conditional independence condition to identify subsets of covariates that are sufficient for identification, implying that the remaining covariates satisfy instrument validity; see, for instance, De Luna et al.
+(2011) and VanderWeele and Shpitser (2011).
+Considering subsets of covariate information may also allow for more efficient causal effect estimation.
+For example, Christgau and Hansen (2024) propose a deep learning-based estimator that learns efficient covariate representations from unstructured data and is asymptotically normal under certain conditions.
+Again, our approach differs from these studies in that it tests the selection-on-observables and instrument validity assumptions jointly, rather than testing the latter conditional on the former.
+Closer to our framework, Entner et al.
+(2013) do not pre-impose the selection-on-observables assumption when searching for a sufficient set of covariates to control for based on conditional independence and consider a parametric modelling approach for testing, however, without providing asymptotic guarantees.
+Here, we suggest a machine learning-based procedure that accommodates nonparametric models and potentially high-dimensional covariates, while having desirable asymptotic properties under certain regularity conditions.
 Our study is also related to several contributions in the computer science-dominated literature on causal discovery, see see e.g.
@@
 The second line of Assumption 1 enforces causal faithfulness, ensuring that only variables which are d-separated in the sense of Pearl (1988), i.e.
-not associated with each other via some causal paths (possibly conditional on other variables) are statistically independent (or conditionally independent).444 d-separation relies on blocking causal paths between variables.
+not associated with each other via some causal paths (possibly conditional on other variables) are statistically independent (or conditionally independent).555d-separation relies on blocking causal paths between variables.
 Formally, a path between two (sets of) variables $A$ and $B$ is blocked when conditioning on a (set of) control variable(s) $C$ if 1.
@@
 Assumption 3 (Conditional dependence between the treatment and instrument).
-[…]
-[…]
+$D\not\!\perp\!\!\!\perp Z|X,$
+where $\not\!\perp\!\!\!\perp$ denotes statistical dependence.
+Together with Assumption 1, which rules out effects of $D$ on $Z$, Assumption 3 either implies that $Z$ causally affects $D$, which is known as first stage effect in the IV literature, or that some (unobserved) characteristics jointly affect $Z$ and $D$ given $X$.
+This assumption is satisfied in Figure 1, where $Z$ affects $D$.
 The fourth assumption invokes selection on observables, i.e. quasi-random treatment assignment conditional on covariates as e.g. considered in Imbens (2004):
@@
 $Y(d){\perp\!\!\!\perp}D|X\quad\forall d\in\mathcal{D},$
-[…]
+where $\perp\!\!\!\perp$} denotes statistical independence.
+Assumption 4 implies that conditional on covariates $X$, there exist no unobserved confounders jointly affecting outcome $Y$ and treatment $D$.
 The fifth assumption invokes instrument validity, requiring that the instrument is conditionally independent of the potential outcomes given the covariates $X$.
@@
 $Y(d){\perp\!\!\!\perp}Z|X\quad\forall d\in\mathcal{D}.$
-[…]
+Assumption 5 rules out unobserved confounders jointly affecting $Y$ and $Z$ when controlling for $X$.
+This is analogous to Assumption 4, but now concerning $Z$ rather than $D$.
+Furthermore, by assuming that the potential outcome is solely a function of $d$ (and not $z$), Assumption 5 also implies that the instrument does not have a direct impact on the outcome, except through its influence on the treatment, conditional on $X$.
+By this exclusion restriction, it holds that conditional on $X$, $Y(d,z)=Y(d,z^{\prime})=Y(d)$ for any instrument values $z$ and $z^{\prime}$.
+Otherwise, Assumption 5 would be violated, because it would follow that $Y(d)=Y(d,Z)$ and $Y(d,Z)\not\!\perp\!\!\!\perp Z|X$.
 The previous assumptions can be used to construct tests for the identification of causal effects.
@@
 For this reason, we subsequently propose a new testing approach that is based on a refined score function that does not require user-selected random terms and entails a test statistic that is normally distributed under both the null and alternative hypotheses.
-[^6] We note that an alternative testing approach that could be applied in our context is suggested by Kook and Lundborg (2024)
+[^7] We note that an alternative testing approach for conditional independence satisfying Neyman orthogonality that can be combined with machine learning and applied in our context is suggested by Lundborg et al.
+(2024), see also Kook and Lundborg (2024).
 To formally discuss the test, let us for the moment assume that the instrument $Z$ is binary and denote by $p(D,X)=\mathbb{P}(Z=1|D,X)$ the instrument propensity score.
@@
 (2006), Hong and White (1995) and Wooldridge (1992).
-In addition to the squared difference in conditional mean outcomes that we already encountered in the previous section, the score function notably contains a term in which the difference in conditional mean outcomes is multiplied with a difference in expressions obtained by inverse probability weighting (IPW) with the instrument propensity score.
+In addition to the squared difference in conditional mean outcomes that we already encountered, the score function notably contains a term in which the difference in conditional mean outcomes is multiplied with a difference in expressions obtained by inverse probability weighting (IPW) with the instrument propensity score.
 In fact, our new score above combines the orthogonalized squared score in (3.3) with the popular doubly robust score, as also considered for testing in Huber and Kueck (2022).
@@
 $+\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))$
-$+\sum_{l=1}^{L}\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}-\theta.$
-This score has a variance that is bounded away from zero, is zero in expectation under the null hypothesis and is Neyman-orthogonal, as formally shown in Appendix A.2.
+$+\sum_{l=1}^{L}\left(\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right).$
+This score has a variance that is bounded away from zero, is zero in expectation under the null hypothesis, $\theta_{0}=0$, and is Neyman-orthogonal, as formally shown in Appendix A.2.
 We may construct cross-fitted estimators of $\theta_{0}$ based on the score function (3.5).
 It is worth noting that the corresponding target parameter in (3.5) is given by
-$\mathbb{E}\left[\sum_{l=1}^{L}(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))^{2}+(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right],$
+$\mathbb{E}\left[\sum_{l=1}^{L}[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))^{2}+(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))]\right],$
 which tests the null hypothesis given in equation (3.1) for binary and discrete instruments, and approximates equation (3.1) in the case of continuous instruments if the bins that define $Z_{l}$ become small.
 When implementing the testing approach, we assume an i.i.d.
-sample of size $n$, in which $i$ denotes the index of an observation and $W_{i}=(Y_{i},D_{i},X_{i},Z_{i})$ are the variable values of observation $i$ in that sample, with $i$in $\{1,2,..,n\}$.
+sample of size $n$, in which $i$ denotes the index of an observation and $W_{i}=(Y_{i},D_{i},X_{i},Z_{i})$ are the variable values of observation $i$ in that sample, with $i$ $\in$ $\{1,2,..,n\}$.
 When testing, we apply cross-fitting as for instance discussed in Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey and Robins (2018) to avoid over-fitting due to a correlation of the estimation of the nuisance parameters and $\theta_{0}$.
@@
 The cross-fitted estimator is given by
-$\hat{\theta}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{N,k}[\psi_{k}(W_{i},0,\hat{\eta})],$ (3.6)
-with $\hat{\eta}=(\hat{\mu},\hat{m},\hat{p}_{1},\dots,\hat{p}_{L})$.
+[…]
+with $\hat{\eta}=(\hat{\mu},\hat{p}_{1},\dots,\hat{p}_{L})$.
 Under the regularity conditions in Assumption 6, the estimator $\hat{\theta}$ is asymptotically normal and $\sqrt{n}$-consistent, as stated in Theorem 1.
@@
 Assumption 6 (Asymptotic Normality).
-The following assumption needs to hold for all $n\geq 3$, $\mathbb{P}\in\mathcal{P}$ and $q>2$: (i) $\|Y\|_{\mathbb{P},q}<C$ and $\mathbb{E}[\varepsilon^{2}1(Z\in Z_{l})]>c$ (ii) Given a random subset $I$ of $[n]$ of size $N=n/K$, the nuisance parameter estimator $\hat{\eta}_{0}=\hat{\eta}_{0}((W_{i})_{i\in I^{c}})$ obeys $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2q}\leq C$, $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},4}\leq\delta_{N}$, and $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2}\leq\delta_{N}^{1/2}N^{-1/4}$ with $\mathbb{P}$-probability not less than $1-o(1)$.
+Define $U=Y-\mu(D,X,Z)$.
+The following assumption needs to hold for all $n\geq 3$, $\mathbb{P}\in\mathcal{P}$ and $q>2$: (i) $\|Y\|_{\mathbb{P},q}<C$ and $\mathbb{E}[U^{2}1(Z\in Z_{l})]>c$ (ii) Given a random subset $I$ of $[n]$ of size $N=n/K$, the nuisance parameter estimator $\hat{\eta}_{0}=\hat{\eta}_{0}((W_{i})_{i\in I^{c}})$ obeys $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2q}\leq C$, $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},4}\leq\delta_{N}$, and $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2}\leq\delta_{N}^{1/2}N^{-1/4}$ with $\mathbb{P}$-probability not less than 1o(1)1-o(1).
 Theorem 1.
 Conditional on Assumptions 6, the estimator in equation (3.6) satisfies
-$\sqrt{n}\sigma^{-1}\hat{\theta}\rightsquigarrow N(\theta_{0},1),$ (3.7)
+$\sqrt{n}\sigma^{-1}\hat{\theta}\overset{d}{\rightarrow}N(\theta_{0},1),$ (3.7)
 uniformly over $P\in\mathcal{P}$, where $\sigma^{2}=E[\psi(W,\theta_{0},\eta_{0})^{2}]$.
@@
 We note that the first step is required to select candidate instruments which, if ultimately selected as instrument $Z$, are statistically significant in terms of satisfying the first stage condition $E[D|X,Z]\neq E[D|X]$.
-Let us assume that $Q$ is low dimensional in the sense that the sample size $n$ is larger than the number of variables in $Q$, denoted by $p$, such that $n>p$.
+Let us assume that $Q$ is low-dimensional in the sense that the sample size $n$ is larger than the number of variables in $Q$, denoted by $p$, $n>p$.
 In this case, step one might be implemented based on a first stage regression of $D$ on $Q$ and selecting all regressors with statistically significant associations after controlling for multiple hypothesis testing issues into the set of candidate instruments $\hat{S}$.
-As statistical criterion for selecting candidate instruments, we suggest first-stage hard thresholding, as performed in the IV selection literature for instance by Guo et al.
-(2018) and Masten and Poirier (2021).
-To formalize the discussion, let $F_{r}$ be the first stage F-statistic in the estimation of $E[D|X,Z]$ when defining $Z=Q_{r}$ and $X=Q_{[r]}$, with $Q_{[r]}=Q\setminus Q_{r}$ (and $Q=X\cup Z$).
-We obtain the estimated set of strong candidate instruments, denoted by $\hat{S}$, as those whose F-statistic exceeds a certain critical value $C_{n}$:
-$\hat{\mathcal{S}}=\{r:F_{r}\geq C_{n}\}$ (4.1)
+In high-dimensional settings where $p>n$, lasso based approaches can be applied to select strong IV candidates.
+Details are given in the next subsection.
+4.1 Strong IV Selection
+To select strong instrumental variables (IVs) in the high-dimensional setting, we use first-stage hard-thresholding (FSHT) proposed by Guo et al.
+(2018).
+In this approach, instruments are considered irrelevant if their t-statistic does not exceed a predefined threshold.
+[^8] Specifically, an instrument is treated as weak if $|t_{\gamma_{j}}|=\left|\frac{\hat{\gamma}_{j}}{\sqrt{\widehat{Var}(\hat{\gamma}_{j})}}\right|<\sqrt{2.01\log(\max(p,n))}.$ (4.1) Here, $\widehat{Var}(\hat{\gamma}_{j})$ denotes the variance estimator of the first-stage coefficient estimator $\hat{\gamma}_{j}$.
+The right-hand side accounts for multiple testing, where the factor $\log(\max(p,n))$ originates from the tail bound of the normal distribution.
+In the low-dimensional case, this term simplifies to $\sqrt{2.01\log(n)}$.
+This procedure is shown to consistently select relevant instruments, as demonstrated in Lemma 1 in Guo et al.
+(2018).
+Windmeijer et al.
+(2021) treat the factor 2.01 as a tuning parameter and show in simulations that this choice of the tuning parameter performs well in practice.
+Meanwhile, Guo et al.
+(2018) argue that the consistent selection property holds as long as the argument inside the logarithm grows with $n$.
+We iteratively consider each variable in $Q$ as instrument $Z$ and any remaining variables as covariates.
+That is, when considering the $j$th variable in $Q$ as instrument and denoting it by $Q_{j}$, we have that $Z=Q_{j}$ and $X=Q_{[j]}$, where $Q_{[j]}=Q\setminus Q_{j}$ and $Q=X\cup Z$ when estimating the first-stage association $E[D|X,Z]$.
+Leveraging the double machine learning (DML) literature, we consider the following partially linear regression (PLR) model specification to estimate the effect of a variable $Z$ on treatment $D$:
+$D$ $=\gamma_{j}^{T}Z+g(X)+\varepsilon,$ (4.2)
+$Z$ $=h(X)+v.$ (4.3)
+Here, $g(X)$ and $h(X)$ are general functions of $X$, and the first-stage effect $\gamma$ of $Z$ is indexed by $j$ to emphasize that it may vary with each variable considered as an instrument.
+Note that although the partial linear model assumes additive separability and a linear effect of $Z$ on $D$, estimation based on this model may also have power in fully nonparametric settings, where the effect of $Z$ on $D$ may be nonlinear and may interact with $X$.
+In such cases, it follows from Stein’s Lemma (Stein, 1981) that $\gamma_{j}$ corresponds to a weighted average of the marginal effects of $Z$ on $D$ given $X$, with all weights being nonzero.
+[^9] In special cases (in particular $Z$ being Gaussian conditional $X$), $\gamma_{j}$ is the (unweighted) average of marginal effects.
+DML-based estimation of $\gamma_{j}$ is asymptotically normal under certain regularity conditions (in particular, when estimators of the models of $D$ and $Z$ attain a convergence rate of $o(n^{-1/4})$), see Chernozhukov et al.
+(2018): $\sqrt{n}\hat{\sigma}_{\gamma,j}^{-1}(\hat{\gamma}_{j}-\gamma_{0,j})\overset{d}{\rightarrow}N(0,1)$, where $\hat{\sigma}_{\gamma,j}^{-1}$ is the standard error of $\hat{\gamma}_{j}$.
+As only one IV is considered in turn, the first-stage F-statistic is equal to the square of the t-statistic, $t_{j}^{2}=F_{j}$, which is asymptotically $\chi_{1}^{2}$-distributed with one degree of freedom, $F_{j}\overset{d}{\rightarrow}\chi_{1}^{2}$ (see Proposition 3 of Masten and Poirier (2021)).
+Under the alternative hypothesis, $\gamma_{j}\neq 0$, we have $\frac{F_{j}}{n}\overset{P}{\rightarrow}\kappa_{j}$, where $\kappa_{j}>0$ is a constant.
+The critical values $C_{n}$ for the F-statistic need to satisfy $C_{n}\to\infty$ and $C_{n}=o(n)$ as $n\to\infty$, for the first-stage hard-thresholding procedure to successfully identify strong candidate instruments with probability approaching 1, i.e.,
+$\lim\limits_{n\rightarrow\infty}P(\hat{S}=\mathcal{S})=1,$
 […]
-An alternative strategy is to iteratively consider each variable in $Q$ as instrument $Z$ and any remaining variables as covariates $X$ when estimating the first stage association $E[D|X,Z]$ based on methods for treatment or policy evaluation like doubly robust estimation, see e.g.
-Robins et al.
-(1994), Robins et al.
-(1995) and Chernozhukov et al.
-(2018).
-Such doubly robust approaches are also applicable in high-dimensional contexts with $n<<p$ when combining them with machine learning to control for important control variables among high-dimensional covariates $X$.
-For instance, the double machine learning algorithm suggested by Soleymani et al.
-(2022), selects direct causal parents in a data-driven manner by iteratively changing the definition of variables whose causal effect is of interest and the control variables.
-It may be applied to select the set of candidate instruments $S$ which are those variables in $Q$ strongly associated with $D$ conditional on the remaining elements in $Q$.
-Second, we proceed to test the null hypothesis in equation (3.2) iteratively over all candidate instruments that pass the first-stage threshold, $\hat{\mathcal{S}}$.
-Our aim is to find a partition of variables for which the conditional independence of the respective candidate instrument holds. To discuss this more formally, we introduce the partition
+4.2 Valid IV Selection and Identification Test
+In the second step, we test the null hypothesis in equation (3.2) iteratively over all candidate instruments that pass the first-stage threshold, $\hat{\mathcal{S}}$.
+Our aim is to find a partition of variables for which the conditional independence of the respective candidate instrument holds.
+To discuss this more formally, we introduce the partition
 $\mathcal{P}_{j}=\{Z=Q_{j},\ \ X=Q_{[j]}\},$
 such that variable $j$ in set $Q$ is chosen to be the instrument, while the remaining variables in set $Q$ are used as control variables.
-Furthermore, let $\mathcal{V}$ denote the set of candidate instruments which are conditionally mean independent of the outcome, satisfying
+Furthermore, let $\mathcal{V}$ denote the set of candidate instruments which are conditionally mean independent of the outcome, thus satisfying condition (2.2) and the null hypothesis in equation (3.2):
 $\mathcal{V}=\{j:E[Y|D,X]=E[Y|D,X,Z]\}.$
-Furthermore, we denote the set of partitions for which the instrument has a first stage effect on the treatment and the conditional independence of the instrument holds by $\mathcal{P}^{*}$:
+Moreover, we denote the set of partitions for which the instrument has a first stage effect on the treatment and the conditional independence of the instrument holds by $\mathcal{P}^{*}$:
 $\mathcal{P}^{*}=\{\mathcal{P}_{j}:j\in\left(\mathcal{S}\cap\mathcal{V}\right)\}.$ (4.4)
 The corresponding estimated set of partition(s) $\mathcal{P}^{*}$, denoted by $\hat{\mathcal{P}}_{pass}$, is given by
-$\hat{\mathcal{P}}_{pass}=\left\{\mathcal{P}_{j}:j\in\left(\hat{\mathcal{S}}\cap\hat{\mathcal{V}}\right)\right\},$ (4.3)
+[…]
 where $\hat{\mathcal{V}}=\{j:|\sqrt{n}\hat{\sigma}^{-1}_{j}\hat{\theta}_{j}|<c_{\alpha}\}$ is the set of instruments for which conditional independence is not rejected based on the test defined in Theorem 1 and Corollary 1 with significance level $\alpha$ and the corresponding critical value $c_{\alpha}$.
 If $|\hat{\mathcal{P}}_{pass}|>1$ such that there is more than one single candidate instrument for which the first stage is sufficiently strong while conditional independence is not rejected, we can proceed in two ways.
-For a set of partitions $\mathcal{P}_{pass}=\{\mathcal{P}_{1},...,\mathcal{P}_{|\mathcal{P}_{pass}|},\}$ where each partition is associated with a specific IV, we define the set of IVs associated with that set of partitions as
-=$\mathcal{I}=\left\{j:\mathcal{P}_{j}\in\mathcal{P}_{pass}\right\},$ (4.4)
+For a set of partitions $\hat{\mathcal{P}}_{pass}=\{\mathcal{P}_{1},...,\mathcal{P}_{|\hat{\mathcal{P}}_{pass}|},\}$ where each partition is associated with a specific IV, we define the set of IVs associated with that set of partitions as
+$\hat{\mathcal{I}}=\left\{j:\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}\right\},$ (4.6)
 we can then select the final partition into IVs and controls as
-$\mathcal{P}_{all}=\{Z=Q_{\mathcal{I}},\ \ X=Q\setminus Q_{\mathcal{I}}\}.$ (4.5)
-Alternatively, we can select the final partition as the one which maximizes the p-value when testing conditional independence:
-$\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\hat{\mathcal{P}}_{pass}),$ (4.6)
-where $p(\mathcal{P})$ is the $p$-value obtained in the conditional independence test of the instrument and the outcome.
+$\hat{\mathcal{P}}_{all}=\{Z=Q_{\hat{\mathcal{I}}},\ \ X=Q\setminus Q_{\hat{\mathcal{I}}}\}$. (4.7)
+In this paper, we select the final partition as the one which maximizes the p-value when testing conditional independence:
+$\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P}_{j}),$ (4.8)
+where $\texttt{p}(\mathcal{P}_{j})$ is the p-value obtained in the conditional independence test of the instrument and the outcome.
 Even more crucial than the question how to choose the final partition if $|\hat{\mathcal{P}}_{pass}|>1$ is the question whether $\hat{\mathcal{P}}_{pass}=\emptyset$ or not.
@@
 To this end, we consider the following hypotheses:
-$H_{0}:\text{no identification}\quad v.s\quad H_{1}:\text{identification}\quad(Y(d){\perp\!\!\!\perp}D|X\quad\forall d\in\mathcal{D}).$ (4.7)
-We conclude that $H_{1}$ is true, i.e, our causal effect of interest is identified if $\hat{\mathcal{P}}_{pass}\neq\emptyset$.
+$H_{0}:\text{no identification}\quad vs.\quad H_{1}:\text{identification (conditional mean independence}).$ (4.9)
+We conclude that $H_{1}$ is true (the ATE is identified, see the null hypothesis in (3.1)) if $\hat{\mathcal{P}}_{pass}\neq\emptyset$.
 The following Theorem 2 helps us to understand the type 1 error of our proposed test.
@@
 If $\hat{\mathcal{P}}_{pass}=\emptyset$, our statement holds trivially.
-Hence, consider a variable $j\in\hat{\mathcal{P}}_{pass}$, i.e, $Z_{j}\in\hat{\mathcal{S}}$ and $Z_{j}\in\hat{\mathcal{V}}$.
+Hence, consider a variable $j:\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}$, i.e, $Z_{j}\in\hat{\mathcal{S}}$ and $Z_{j}\in\hat{\mathcal{V}}$.
 Therefore, it holds
@@
 3:  Select strong instruments $\hat{\mathcal{S}}$:
-4:     Regress: D $\leftarrow$ Q
-5:     $\hat{F}_{j}\leftarrow\hat{t}_{j}^{2}$
-6:     if $\hat{F}_{j}>C_{0.1/log(n)}$ then $j\in\hat{\mathcal{S}}$
-7:     Select instruments passing conditional independence test $\hat{\mathcal{V}}$:
-8:     for $j\in\hat{\mathcal{S}}$ do
-9:         $\hat{\mu}(D,X,Z)\leftarrow$ and $\hat{m}(D,X)\leftarrow$ and $\hat{p}(D,X)\leftarrow$ estimates from machine learning
-10:         $\hat{\theta}\leftarrow$ constructed via score in equation (3.4) (binary case) or (3.5) (continuous case)
-11:         $\hat{t}_{ind,j}\leftarrow$ test $H_{0}:\hat{\theta}=0$
-12:         if $\hat{t}_{ind,j}<t_{crit}$ then $j\in\hat{\mathcal{V}}$
-13:     Select final partition $\hat{\mathcal{P}}$:
-14:     $\hat{\mathcal{P}}_{pass}\leftarrow\{j\in\hat{\mathcal{S}},j\in\hat{\mathcal{V}}\}$
-15:     if $\hat{\mathcal{P}}_{pass}=\emptyset$ then end and report that identification is “rejected”
-16:     else if $|\hat{\mathcal{P}}_{pass}|=1$ then $\hat{\mathcal{P}}\leftarrow\hat{\mathcal{P}}_{pass}$
-17:     else if $|\hat{\mathcal{P}}_{pass}|>1$ then
-18:         $p(\hat{\mathcal{P}}_{pass})\leftarrow\text{p-value}(\hat{t}_{ind,j})$ for $j\in\hat{\mathcal{V}}$
-19:         $\hat{\mathcal{P}}=\underset{\mathcal{P}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P})$
-20:     In case identification is confirmed ($\hat{\mathcal{P}}_{pass}\neq\emptyset$), estimate the treatment effect:
-21:     Regress: $Y\leftarrow D$ and $X=Q_{[j]}\in\hat{\mathcal{P}}$
+4:  for $j\in Q$ do
+5:   Predict $D$ by $Q_{j}$ conditional on $Q\setminus Q_{j}$ using machine learning or regression
+6:   $\hat{F}_{j}\leftarrow\hat{t}_{j}^{2}$
+7:   if $\hat{F}_{j}>C_{\tilde{\alpha}}$ with $\tilde{\alpha}=0.1/log(n)$ $j\in\hat{\mathcal{S}}$ then
+8:  Select instruments passing conditional independence test $\hat{\mathcal{V}}$:
+9:  for $j\in\hat{\mathcal{S}}$ do
+10:   $\hat{\mu}(D,X,Z)\leftarrow$ and $\hat{p}(D,X)\leftarrow$ estimates from machine learning
+11:   $\hat{\theta}_{j}\leftarrow$ constructed via score in equation (3.4) (binary case) or (3.5) (continuous case)
+12:   $\hat{\sigma}_{j}\leftarrow$ constructed as described in Theorem 1
+13:   $\hat{t}_{ind,j}=\frac{\hat{\theta}_{j}}{\hat{\sigma}_{j}}\leftarrow$ test $H_{0}:\theta=0$
+14:   if $\hat{t}_{ind,j}<c_{\alpha}$ then $j\in\hat{\mathcal{V}}$
+15:  Select final partition $\hat{\mathcal{P}}$:
+[…]
+17:  if $\hat{\mathcal{P}}_{pass}=\emptyset$ then end and report that $H_{0}$ is not rejected (no identification)
+18:  else if $|\hat{\mathcal{P}}_{pass}|=1$ then $\hat{\mathcal{P}}\leftarrow\hat{\mathcal{P}}_{pass}$
+19:  else if $|\hat{\mathcal{P}}_{pass}|>1$ then
+20:   $\texttt{p}(\mathcal{P}_{j})\leftarrow\texttt{p}-value(\hat{t}_{ind,j})$ for $j\in\hat{\mathcal{V}}$
+21:   $\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P}_{j})$
+22:  In case identification is confirmed ($\hat{\mathcal{P}}_{pass}\neq\emptyset$), estimate the treatment effect:
+23:  Regress: $Y\leftarrow D$ and $X=Q_{[j]}\in\hat{\mathcal{P}}_{pmax}$
 5 Simulation study
@@
 $D$ $=$ $I\{X^{\prime}\beta+Z+\delta W+V>0\},$
-$X,Z$ $\sim$ $binomial(\pi),$
+$X,Z$ $\sim$ $Bernoulli(\pi),$
 $W$ $\sim$ $N(0,1),U\sim N(0,1),V\sim N(0,1),$
@@
 While the supposed instrument $Z$ is binary, the unobserved terms $U,V,W$ are random, standard normally distributed variables that are independent of each other, of $Z$, and of $X$.
-Covariates $X$ are created as follows.
-We first draw a matrix of multivariate normal variables $\tilde{X}\sim N(\mathbf{0},\bm{\Sigma})$ where $\bm{\Sigma}$ is a covariance matrix with $Cov(X_{j},X_{k})=0.5^{|j-k|}$.
-We then compute probabilities $\pi_{i}=\frac{1}{1+exp(-\tilde{X}_{i})}$ and draw $X_{i}\sim Bernoulli(\pi_{i})$.
-$\beta$ gauges the effects of the covariates on $Y$ and $D$, respectively, and thus, the magnitude of confounding due to observables.
-The $j$th element in the coefficient vector $\beta$ is set to $0.8/j$ for $j=1,...,4$, implying a linear decay of covariate importance in terms of confounding for the first 4 covariates.
-We use the final element, $j=20$ as the instrument $Z$.
-This results in 20 total covariates, with the first 4 as confounders, and the last as the instrument $Z$.
-We investigate the performance of our testing approach in $100$ simulations with sample sizes of $n=1000$, $4000$, and $16000$, setting the number of covariates to $20$.
-Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with two folds ($K=2$) when using the doubly robust score function in equation (3.4).
-For nuisance parameter estimation, lasso regression with default parameters as implemented in the glmnet package by Friedman et al.
+The joint set of covariates and instruments $Q$ is created as follows.
+We first draw a matrix of multivariate normal variables $\tilde{Q}\sim N(\mathbf{0},\bm{\Sigma})$ where $\bm{\Sigma}$ is a covariance matrix with $Cov(\tilde{Q}_{j},\tilde{Q}_{k})=0.5^{|j-k|}$.
+We then compute probabilities $\pi_{i}=\frac{1}{1+exp(-2\cdot\tilde{Q}_{i})}$ and draw $Q_{i}\sim Bernoulli(\pi_{i})$.
+There are in total 10 variables in the set of observed variables $Q=(X,Z)$, with the last one being the (single) instrument $Z$.
+Concerning the first nine elements in $Q$, which are considered as covariates $X$, $\beta$ gauges their effects on $Y$ and $D$, respectively, and thus, the magnitude of confounding due to observables.
+The $j$th element in the coefficient vector $\beta$ is set to $0.8/j$ for j=1,,4j=1,...,4, implying a linear decay of covariate importance in terms of confounding for the first 4 covariates.
+Furthermore, for elements j=5,,9j=5,...,9, the coefficients are equal to zero, implying that covariates 5 to 9 are random noise variables that are neither confounders, nor instruments.
+In a second simulation design, we also consider continuous variables $Q$, based on the following data generating process:
+$Y$ $=$ $D+X^{\prime}\beta+\gamma Z+W+U,$
+$D$ $=$ $I\{X^{\prime}\beta+Z+\delta W+V>0\},$
+$X,Z$ $\sim$ $Uniform(-0.5,0.5),$
+$W$ $\sim$ $N(0,1),U\sim N(0,1),V\sim N(0,1).$
+The set of observed covariates $X$ and the instrument $Z$ are now drawn from the multivariate uniform distribution, $X,Z\sim Uniform(-0.5,0.5)$.
+All other properties of this data generating process are identical to the previous scenario with binary $Z$ and $X$ variables.
+Specifically, the outcome $Y$ is characterized by the same linear function of $D,X,W$, and $U$ (and the supposed instrument $Z$ in the case where $\gamma\neq 0$) as before, and we again consider 10 variables in the set $Q=(X,Z)$, with the last one being the (single) instrument $Z$, elements j=1,,4j=1,...,4 being observed confounders, and elements j=5,,9j=5,...,9 being random noise variables that are neither confounders nor instruments.
+We investigate the performance of our testing approach in $200$ simulations with sample sizes of $n=1000$, 40004000, and $16000$.
+For estimating the (conditional) first stage effect of elements in $Q$ on $D$ to select sufficiently strong candidate instruments, we use Double Machine Learning (DML) for partially linear models (Chernozhukov et al., 2018).
+DML can be applied to both discrete and continuous elements in $Q$ and is implemented in the DoubleML package for statistical software R.
+The so-called nuisance parameters in the first stage are estimated using lasso regression with cross-fitting, setting the number of folds to five (K = 5).
+Specifically, the classification lasso is used for discrete elements in $Q$, modelling probabilities for treatment assignment while lasso regression is used for continuous variables.
+For testing the conditional independence of any selected candidate instrument in the second stage, we consider a cross-fitted estimator $\hat{\theta}$ of equation (3.6) with five folds ($K=5$), which is based on the doubly robust score functions in equations (3.4) and (3.5) for binary and continuous candidate instruments, respectively.
+In the case of a continuous candidate instrument, its support is partitioned based on the quartiles of its distribution, such that $L=4$.
+For nuisance parameter estimation in the second stage, lasso regression with default parameters as implemented in the glmnet package by Friedman et al.
 (2010) for the statistical software R is applied.
-We consider several statistics from our simulations: the estimated violation, $\hat{\theta}$, when using variable $Z$ as the instrument (est), the standard deviation of these estimates (std), and the average of the standard errors across all samples (mean se).
-These statistics are useful for judging the performance of the test conditional on selecting $Z$ as the instrument, which is the correct choice if $\gamma=0$, while there is no valid instrument if $\gamma\neq 0$.
-However, since the choice of the instrument is not fixed a priori but is part of the estimation process, we also consider the following two performance measures.
-The first one is the empirical detection rate of $Z$ (det.Z), which indicates the frequency with which variable $Z$ is selected as the best candidate instrument and has a p-value greater than 10% when testing conditional independence based on the estimate $\hat{\theta}$.
-In other words, this corresponds to the proportion of simulations in which our method simultaneously selects $Z$ as the instrument and keeps the null hypothesis.
-In scenarios where the null hypothesis holds ($\delta=\gamma=0$), this detection rate should approach one as the sample size increases.
-The second measure is an equivalent detection rate for the confounders $X_{1}$ to $X_{4}$ (det.X).
-It indicates the frequency with which one of the confounders is selected as the best candidate instrument and yields a p-value greater than 10% when testing conditional independence using this confounder as the instrument.
-As the sample size increases, this detection rate should always approach zero, because confounders are never valid instruments.
-Table 1: Simulations: Binary Instrument
-N est std mean se det.Z det.X
+We consider several statistics from our simulations: the estimated violation, $\hat{\theta}$, when using variable $Z$ as the instrument, the standard deviation of the estimated violation (std), and the average of the standard error of the estimated violation across all simulation samples (mean se).
+These statistics are useful for judging the performance of the test conditional on selecting $Z$ as the instrument, which is the correct choice if $\gamma=0$, while there is no valid instrument if $\gamma\neq 0$ and no identification when controlling for covariates if $\delta\neq 0$.
+Since the choice of the instrument is not fixed a priori but is part of the estimation process, we also consider three further statistics.
+The first one is the empirical selection rate of $Z$ (sel.Z), which indicates the frequency with which variable $Z$ is selected as the best candidate instrument and also has a p-value greater than 30% when testing conditional independence based on the estimate $\hat{\theta}$.
+[^10] We suggest the threshold of 30% based on examination of the distribution of p-values for both true instruments and confounders, see Figures A1 and A2.
+This stricter threshold guards against the possible choice of a confounder as an instrument, although a more traditional threshold of 10% could also be used.
+In other words, this corresponds to the proportion of simulations in which our method simultaneously selects $Z$ as the instrument and at the same time keeps the null hypothesis of conditional independence.
+In scenarios where the null hypothesis holds ($\delta=\gamma=0$), this selection rate should approach one as the sample size increases.
+The second statistic is the analogous selection rate of the noise variables $X_{5}$ to $X_{9}$ (sel.noconf).
+It indicates the share of simulations in which one of the non-confounders (that are not instruments either) is selected as the best candidate instrument and at the same time yields a p-value greater than 30% when testing conditional independence when using this noise variable as the instrument.
+As the sample size increases, this selection rate should always approach zero, because asymptotically, noise variables do not have a first stage effect on the treatment conditional on other elements in $Q$.
+Therefore, noise variables should not be selected as candidate instrument for the conditional independence test.
+Finally, we report an equivalent selection rate for the confounders $X_{1}$ to $X_{4}$ (sel.conf).
+It corresponds to the share of simulations in which one of the confounders is selected as the best candidate instrument and at the same time yields a p-value greater than 30% when testing conditional independence when using this confounder as instrument.
+As the sample size increases, this selection rate should always approach zero, because confounders are never valid instruments.
+Table 1: Simulations: Single Instrument
+PANEL A: Binary Instrument
+N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
 Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
-1000 0.002 0.039 0.018 31% 48%
-4000 -0.001 0.010 0.006 46% 27%
-16000 -0.000 0.003 0.002 61% 0%
+1000 32% 2% 59% 0.002 0.010 0.008
+4000 45% 1% 42% 0.000 0.002 0.002
+16000 68% 2% 4% 0.000 0.000 0.000
 Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
-1000 -0.027 0.029 0.019 39% 21%
-4000 -0.027 0.014 0.010 8% 21%
-16000 -0.026 0.006 0.006 0% 0%
+1000 50% 3% 32% 0.012 0.012 0.016
+4000 15% 4% 44% 0.012 0.006 0.009
+16000 0% 4% 20% 0.013 0.003 0.005
 Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$)
-1000 -0.353 0.185 0.146 6% 69%
-4000 -0.457 0.111 0.091 0% 72%
-16000 -0.578 0.072 0.053 0% 2%
-Notes: columns ‘est’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z$ as instrument, its standard deviation, and the average of the standard errors across all samples, respectively.
-‘det.Z’ gives the empirical detection rate of selecting $Z$ as instrument and obtaining a p-value $>0.10$ (or 10%) when estimating $\theta$.
-This corresponds to the frequency with which $Z$ is selected as the best candidate instrument and also entails a p-value of more than 10% when testing conditional independence.
-‘det.X’ is an equivalent measure for confounders $X_{1}$ to $X_{4}$.
-It corresponds to the frequency with which one of the confounders is selected as the best candidate instrument and also entails a p-value of more than 10% when testing conditional independence.
-Table 1 reports the simulation results for a binary $Z$ and binary $X_{j}$.
-$Z$ is correlated with $X_{j}$ according to the covariance matrix $\bm{\Sigma}$.
-The top rows focus on the case where $\delta=\gamma=0$, meaning that both Assumptions 4 and 5 are satisfied.
-As the sample size increases, the test is more likely to simultaneously select the true instrument $Z$ and keep the null hypothesis that the assumptions are satisfied.
-Specifically, the detection rate of $Z$ (det.Z) increases from 31% with the smaller sample size of 1000 observations to 61% with 16000 observations.
-This improvement is mirrored by a reduction in the rate at which a confounder variable ($X_{1}...X_{4}$) is selected as the instrument and passes the test (det.X).
-Under the largest sample size of 16000 observations, it never occurs that any confounder is selected as the best instrument and passes the conditional independence test.
-Additionally, we observe that, conditional on having selected the true instrument $Z$, the estimated violation $\hat{\theta}$ (est) is close to zero for any sample size.
-The intermediate rows present the results for a violation of Assumption 4, selection on observables, when setting $\delta=2$, $\gamma=0$.
-As the sample size increases, the detection rates of $Z$ (det.Z) and the confounder variables (det.X) both go to zero, as expected.
-Furthermore, the estimated violation $\hat{\theta}$ conditional on using $Z$ as an instrument (est) is now bounded away from zero across all samples.
-The lower rows of Table 1 provide the performance measures under a violation of Assumption 5, instrument validity, when considering $\delta=0$, $\gamma=0.5$.
-Again and as it should be the case, both the detection rates of $Z$ (det.Z) and the confounder variables (det.X) tend to zero as the sample size increases.
-At the same time, the estimated violation conditional on using $Z$ as an instrument (est) is substantially biased across all samples.
+1000 30% 2% 57% 0.045 0.040 0.050
+4000 3% 2% 68% 0.053 0.019 0.027
+16000 0% 3% 8% 0.054 0.010 0.014
+PANEL B: Continuous Instrument
+N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
+Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
+1000 55% 2% 19% -0.007 0.029 0.031
+4000 65% 4% 8% -0.001 0.015 0.016
+16000 64% 1% 0% 0.000 0.007 0.009
+Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
+1000 16% 6% 6% -0.018 0.031 0.017
+4000 8% 2% 1% -0.015 0.018 0.007
+16000 0% 2% 0% -0.030 0.014 0.003
+Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$)
+1000 0% 4% 18% -0.490 0.192 0.049
+4000 0% 6% 8% -0.552 0.103 0.026
+16000 0% 1% 0% -0.554 0.055 0.013
+Notes: columns $\hat{\theta}$’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z$ as instrument, its standard deviation, and the average of the standard errors across all samples, respectively.
+‘sel.Z’ gives the share of simulations in which $Z$ is selected as best candidate instrument and simultaneously yields a p-value >0.30>0.30 (or 30%) when testing conditional independence based on the estimate of $\hat{\theta}$.
+‘sel.noconf’ is an equivalent measure for covariates $X_{5}$ to $X_{9}$, corresponding to the share of selecting a non-confounder as best candidate instrument and having a p-value >> 30% when testing conditional independence.
+‘sel.conf’ is an equivalent measure for covariates $X_{1}$ to $X_{4}$, corresponding to the share of selecting an observable confounder the best candidate instrument and having a p-value >> 30% when testing conditional independence.
+Figure 2: Simulations: Distribution of Estimates
+This figure depicts density plots of the estimated violation $\hat{\theta}$ for the true IV (covariate $X_{10}$), across the simulation settings.
+The top and bottom rows are the binary and continuous IV cases, respectively.
+The first plots on the left provide results for $\delta=\gamma=0$, i.e.
+Assumptions 4 and 5 are satisfied.
+The middle plots depict the scenario of Assumption 4 being violated, and the right-most plots Assumption 5 is violated.
+Orange is the smallest sample size (N = 1000), green the medium sample size (N=4000), and purple is the largest (N=16000).
+Table 1 reports the simulation results for binary variables in $Q$ in Panel A, and for continuous elements in $Q$ in Panel B, respectively.
+In both panels, the top rows provide the results for $\delta=\gamma=0$, implying that both Assumptions 4 and 5 are satisfied and conditional independence holds.
+As the sample size increases, the test is more likely to simultaneously select the true instrument $Z$ and keep the null hypothesis of conditional independence.
+Specifically, the selection rate of $Z$ (sel.Z) increases from 32% with the smaller sample size of 1000 observations to 68% with 16000 observations for the binary instrument, and from 55% with 1000 observations to 64% with 16000 observations for the continuous instrument.
+This improvement is mirrored by a reduction in the rate at which a confounder variable ($X_{1}...X_{4}$) is selected as the instrument and passes the test (sel.conf).
+Under the largest sample size, it only occurs that any confounder is selected as the best instrument and passes the conditional independence test 4% of the time for the binary and never for the continuous scenario.
+However, we observe that the noise variables ($X_{5}...X_{9}$) sometimes appear to step in, as their selection rate (sel.noconf) may transiently increase in response to the decreasing selection rate of confounders.
+Additionally, we observe that, conditional on having selected the true instrument $Z$, the estimated violation $\hat{\theta}$ is close to zero for any sample size and never statistically significant at conventional significance levels.
+Figure 2 depicts the distribution of the estimated $\hat{\theta}$ and visually confirms this finding in the first column.
+The top row of density plots show the binary $\hat{\theta}$ and the bottom row shows the continuous estimates.
+As sample size increases, the distribution moves closer to zero, as expected.
+The intermediate rows of Table 1 and middle column of Figure 2 present the results for a violation of Assumption 4, selection on observables, when setting $\delta=2$, $\gamma=0$.
+As the sample size increases, the selection rates of $Z$ (sel.Z) and the confounder variables (sel.conf) both go towards zero, as expected.
+The rate of finding a nonempty set $\hat{\mathcal{P}}_{pass}$, which is equivalent the sum of the selection rates sel.Z, sel.noconf and sel.conf, also goes towards zero, finding no identification as expected.
+Furthermore, the estimated violation conditional on using $Z$ as an instrument ($\hat{\theta}$) is statistically different from zero under the larger sample sizes $n=4000,16000$.
+We note that in settings where Assumption 4 is violated, the selection rate of a confounder (sel.conf) in the binary scenario is still slightly high, around 20% in the largest sample for the binary instrument, athough this rate is 0% for the continuous instrument.
+[^11] In further experiments (not reported here), strengthening the degree of confounding by increasing the value of $\gamma$ leads to a substantial reduction in the rate at which a confounder is selected as the best candidate instrument.
+The lower rows of Table 1 and final column of Figure 2 provide the performance measures under a violation of Assumption 5, instrument validity, when considering $\delta=0$, $\gamma=0.5$.
+Again and as it should be the case, both the selection rates of $Z$ (sel.Z) and the confounder variables (sel.conf) tend to zero as the sample size increases.
+At the same time, the estimated violation conditional on using $Z$ as an instrument ($\hat{\theta}$) is large in absolute terms and highly statistically significant across all sample sizes, and $\hat{\mathcal{P}}_{pass}$ is empty 89% of the time for binary and 99% of the time for continuous instruments for the largest sample size.
+We now consider a modification of the previous data generating process that entails the existence of multiple instruments.
+Specifically, the true set of instruments $(Z_{1}...Z_{3})$ now comprises the last three elements in $Q$, i.e.
+variables 8 to 10, which are constructed equivalently to $Z$ in the previous setup.
+We investigate the performance when Assumption 5 is violated for all of the instruments.
+This implies that the set of confounders remains the same as before ($X_{1}...X_{4}$), while the set of noise variables now only consists of three elements ($X_{5}...X_{7}$).
+Results for the multiple instrument simulations are shown in Table 2, and are similar to those shown in the single instrument results.
+In the case of multiple binary instruments and no violations of the assumptions, the selection rate of $Z$ (sel.Z) is up to 96% in the largest sample size.
+As in the single instrument scenario, under a violation of Assumption 4, selection on observables, the selection rates of the true instruments (sel.Z) and confounding variables (sel.conf) go to zero.
+For multiple continuous instruments (Panel B), results are consistent with the single instrument setting, with the exception of the slightly lower increase of the instrument selection rate in the larger sample size compared to the binary version (67% in the case of multiple instruments, compared to 64% under the single instrument scenario).
+However, in all settings of the multiple continuous instrument case, the rate of selecting a confounder (sel.conf) goes to zero and the estimate magnitudes and $\hat{\mathcal{P}}_{pass}$ behave as expected, moving towards zero and the empty set under no violation of the assumptions and away from zero otherwise.
+Table 2: Simulations: Multiple Instruments
+PANEL A: Binary Instruments
+N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
+Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
+1000 80% 1% 18% 0.002 0.008 0.007
+4000 84% 2% 12% 0.000 0.002 0.002
+16000 96% 0% 0% 0.000 0.000 0.000
+Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
+1000 86% 1% 11% 0.010 0.012 0.014
+4000 40% 2% 40% 0.010 0.005 0.007
+16000 0% 3% 16% 0.011 0.003 0.004
+Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$ $\forall Z$)
+1000 70% 0% 21% 0.053 0.035 0.050
+4000 15% 2% 48% 0.058 0.018 0.026
+16000 0% 1% 8% 0.058 0.009 0.013
+PANEL B: Continuous Instruments
+N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
+Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
+1000 62% 1% 6% -0.010 0.035 0.031
+4000 64% 1% 4% -0.003 0.017 0.016
+16000 67% 2% 0% -0.001 0.010 0.008
+Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
+1000 38% 2% 2% -0.027 0.036 0.017
+4000 14% 0% 1% -0.025 0.021 0.007
+16000 0% 0% 0% -0.040 0.007 0.003
+Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$ $\forall Z$)
+1000 20% 2% 10% -0.116 0.090 0.037
+4000 0% 2% 6% -0.142 0.064 0.019
+16000 0% 2% 0% -0.140 0.017 0.010
+Notes: columns $\hat{\theta}$’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z_{3}$ (the last element in $Q$) as instrument, its standard deviation, and the average of the standard errors across all samples, respectively.
+‘sel.Z’ gives the share of simulations in which one of the instruments $Z_{1}$ to $Z_{3}$ is selected as best candidate instrument and simultaneously yields a p-value >0.30>0.30 (or 30%) when testing conditional independence based on the estimate of $\hat{\theta}$.
+‘sel.noconf’ is an equivalent measure for covariates $X_{5}$ to $X_{7}$, corresponding to the share of selecting a non-confounder as best candidate instrument and having a p-value >> 30% when testing conditional independence.
+‘sel.conf’ is an equivalent measure for covariates $X_{1}$ to $X_{4}$, corresponding to the share of selecting an observable confounder the best candidate instrument and having a p-value >> 30% when testing conditional independence.
 Our simulations point to the consistency of our procedure in terms of selecting the correct instrument when Assumptions 4 and 5 are satisfied, and the rejection of the null hypothesis when violations of these assumptions occur.
-However and as a word of caution for empirical applications, our simulations also show that the sample size might need to be quite substantial for our testing approach to perform adequately, as the finite sample performance in our simulations is not convincing in moderate samples of just several thousand observations.
+Where there is more than one binary candidate instrument and both assumptions are satisfied, the procedure chooses one of the correct instruments $90\%$ of the time (sel.Z) with the largest sample size (N = 16000).
+However and as a word of caution for empirical applications, our simulations also show that the sample size might need to be quite substantial for our testing approach to perform adequately.
 6 Empirical application
-This section provides an empirical application to the Job Corps (JC) training program, one of the largest and most comprehensive job training programs for disadvantaged youth in the US that offers various types of academic and vocational training.
-The data come from the National Job Corps Study, a randomized experiment conducted in the mid-to-late 1990s in the United States to evaluate the effectiveness of JC on various labour market outcomes.
-Schochet et al.
-(2001) and Schochet et al.
-(2008) evaluate the impact of random program assignment on a wide range of labor market outcomes and find positive effects on education, employment, and earnings in the longer run.
-We employ our testing methodology to analyze the experimental JC data as available in the causalweight package for the statistical software R, provided by Bodory and Huber (2018).
-This dataset named JC comprises 9240 observations and 46 variables.
-Among these variables is a binary indicator denoting participation in any form of academic or vocational training within the first year following program assignment, which serves as our treatment variable of interest $D$.
-It is noteworthy that while assignment to JC is randomized, actual participation in training might be selective due to non-compliance with the assignment.
-For instance, some individuals eligible for JC may choose not to participate.
-Our outcome variable are weekly earnings in the fourth year post-JC assignment.
-Alongside random program assignment, which stands as an obvious candidate instrument, our vector $Q$ encompasses 28 pre-assignment characteristics.
-These include gender, age, ethnicity, education level, mother tongue, marital status, a binary indicator for having at least one child, previous labor market participation and earnings, household size, parents’ education, welfare receipt during childhood, as well as variables related to health and health behavior.
-Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with two folds ($K=2$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively.
+This section provides an empirical application to the Oregon Health Insurance Experiment (OHI), in which a group of low-income adults were selected randomly by lottery to be eligible to apply for Medicaid, a type of public health insurance in the United States.
+The experiment began in early 2008, when the state of Oregon drew names from the 90,000 individuals who had signed up for a limited number of spaces in the Medicaid program, and a rolling notification of participants regarding their lottery outcomes ran until October 2009.
+Previous work uses the random program assignment as an instrument for health insurance coverage and finds increases in primary, emergency, and preventive care utilisation, decreases in out-of-pocket medical expenditures, and other improvements in self-reported physical and mental health (Finkelstein et al., 2012; Baicker et al., 2013; Taubman et al., 2014; Finkelstein et al., 2019).
+We apply our testing methodology to the experimental OHI data, which is publicly available from the NBER at https://www.nber.org/research/data/oregon-health-insurance-experiment-data, using the sample definition adopted by Finkelstein et al.
+(2012).
+The data comprises 23,762 observations and 692 variables.
+Among the variables is a binary indicator denoting whether an individual was ever enrolled in Medicaid following rolling notification of lottery selection by October 2009.
+This serves as the treatment variable of interest $D$.
+It is noteworthy that while assignment to OHI is randomized, actual enrolment in Medicaid might be selective due to non-compliance with the assignment.
+For instance, some individuals selected as eligible for Medicaid may choose not to enrol.
+We examine two outcome variables: the number of primary care visits and the number of prescription medications taken.
+Alongside random program assignment, which stands as an obvious candidate instrument, our vector $Q$ encompasses 218 pre-assignment characteristics.
+These include gender, age, ethnicity, education level, household size, employment and labour income, as well as a suite of variables capturing pre-treatment health status, healthcare utilisation and health expenditure.
+We also include individual indicator variables for whether there are any missing values for each pre-treatment characteristic.
+Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with five folds ($K=5$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively.
 In the case of a continuous candidate instrument, its support is partitioned based on the quartiles of its distribution, such that $L=4$.
-As in the simulations, we apply lasso regression for nuisance parameter estimation.
-When running our algorithm, it selects random program assignment as the instrument, yielding a very high p-value of 76% when running the conditional independence test.
+As in the simulations, we apply lasso regression for nuisance parameter estimation, although we also report estimates using random forest and gradient boosting as alternatives.
+Furthermore, we drop observations with extremely low or high propensity scores, which cause estimation instability and lead to high variance.
+For a discussion of such trimming based on the propensity score, see, for instance, Crump et al.
+(2009) and Lechner and Strittmatter (2019).
+Specifically, we discard observations for which the estimated propensity score $p_{l}(D,X)=P(Z\in Z_{l}\mid D,X)$, defined for a partition $Z_{l}$ of the candidate instrument (whether continuous or discrete), falls outside the interval $0.01<p_{l}(D,X)<0.99$.
+We require that no more than 5% of observations be trimmed using this rule in order for the candidate instrument to be considered.
+Table 3: Empirical Application
+PANEL A: Primary Care Visits
+Variable $\hat{\theta}$ se p-value
+Lasso 0.000 0.000 0.997
+Random Forest -0.000 0.013 0.999
+XGBoost -0.070 0.207 0.735
+PANEL B: Number of Prescriptions
+Method $\hat{\theta}$ se p-value
+Lasso 0.001 0.000 0.148
+Random Forest -0.007 0.010 0.472
+XGBoost 0.073 0.088 0.407
+Notes: this table reports the estimate $\hat{\theta}$ from equation (3.6) with five folds ($K=5$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively.
+‘se‘ and ‘p-value‘ report the standard error and reported p-value for estimate $\hat{\theta}$ on the random program assignment variable.
+Results are depicted in Table 3, where we show the estimate $\hat{\theta}$ for the random program assignment variable.
+Although there are multiple candidate instruments in $\hat{S}$, none of the other candidates were retained as possible IVs, as their instrument propensity score is either too large or too small to satisfy the trimming rule.
+[^12] We report every variable selected in the first stage to be included in $\hat{S}$, sorted by decreasing p-value, in Appendix Table A1, as well as the number of observations dropped following the trimming rule.
+When running our algorithm, it selects random program assignment as an instrument, yielding estimates of $\hat{\theta}$ with values very close to zero.
+According to our chosen p-value threshold of 30%, the random assignment is a valid instrument across all three algorithms for the doctor visits outcome.
+In the case of the lasso algorithm, the estimate is exactly zero and a p-value of 99.7% is reported, when running the conditional independence test.
 On the one hand, this indicates the validity of the instrumental variable (IV), implying not only the randomness of program assignment inherent to the experimental design.
-It also suggests that the assignment does not directly affect the earnings outcome other than through the treatment (e.g.
-through motivation or disappointment when being or not being eligible for the program).
+It also suggests that the assignment does not directly affect the earnings outcome other than through the treatment (e.g., through motivation or disappointment when being or not being eligible for the program).
 On the other hand, the testing result suggests that the selection-on-observables assumption holds for the treatment when controlling for the pre-assignment covariates available in the data.
-Consequently, given the covariates, our result suggests that we may evaluate the average treatment effect (ATE) on the total population.
+For the number of prescriptions outcome, the test results are less conclusive: the random forest algorithm yields a p-value of 0.472 and the XGBoost algorithm yields a p-value of 0.407, passing the 0.3 threshold, but the lasso results in a lower p-value.
+Consequently, given the covariates, our result suggests that we may evaluate the average treatment effect (ATE) on the total population for the doctor visits outcome.
 In contrast, employing an IV-based approach utilizing random assignment as the instrument for effect estimation would, under specific additional assumptions like treatment monotonicity in the instrument, only permit assessing the local average treatment effect (LATE) on the subpopulation of compliers, whose participation in training aligns with the random assignment, as discussed in Imbens and Angrist (1994) and Angrist et al.
@@
 In this paper, we introduced a machine learning-based algorithm based on a novel doubly robust score function designed to detect and test, in a data-adaptive manner, the presence of control variables sufficient for identifying treatment effects in observational data, as well as variables satisfying instrument validity.
+The algorithm may conclude that identification is given, enabling the estimation of the ATE as well as LATE or conditional average treatment effects (CATEs).
 The method learns the partition of observed variables into instruments and control variables from the data, relying on the condition that valid instruments are conditionally independent of the outcome given the treatment and control variables that imply the conditional treatment exogeneity.
@@
 However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations.
-Finally, we applied our algorithm to empirical labor market data from the Job Corps study.
-Our approach (as expected) selected random assignment into the Job Corps program as valid instrument, while also suggesting that training participation in the first year after assignment, our treatment of interest, is exogenous conditional on pre-assignment covariates.
+Further experiments (not reported in this work) improved the performance of the test in terms of reducing the rate at which a confounder is chosen as the candidate instrument, when the strength of violations of Assumptions 4 and 5 are increased by increasing the values of $\gamma$ and $\delta$.
+However, as the strength of these violations may not be so severe in practice, we maintain lower values for $\gamma$ and $\delta$ in our reported simulation results.
+Finally, we applied our algorithm to empirical health data from the Oregon Health Insurance Experiment.
+Our approach (as expected) always selected random assignment into the program as valid instrument, across a variety of algorithms for the nuisance function learner.
 Appendix A Appendix
@@
 Equation (3.4) suggests the following score function for testing when $Z$ is binary:
-$\quad\tilde{\psi}(W,\theta,\eta)$ (A.1)
+$\qquad\tilde{\psi}(W,\theta,\eta)$ (A.1)
 $=(\mu(D,X,1)-\mu(D,X,0))^{2}$
@@
 with
-$\quad\tilde{\psi}_{1}(W,\theta,\eta)$
+$\qquad\tilde{\psi}_{1}(W,\theta,\eta)$
 $=(\mu(D,X,1)-\mu(D,X,0))^{2}$
@@
 Neyman orthogonality of $\tilde{\psi}$ can be shown by taking the Gateaux derivates w.r.t. the nuisance parameters:
-$\quad\partial_{r}E[\tilde{\psi}_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
+$\qquad\partial_{r}E[\tilde{\psi}_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big|_{r=0}$
 $=E\left[\partial_{r}\tilde{\psi}_{1},(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big|_{r=0}\right]$
 $=2E\left[((\mu_{0}(D,X,1)-\mu_{0}(D,X,0))((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))\}\right]$
-$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
-$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
-$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(Y-\mu_{0}(D,X,1))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
-$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
-$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
-$\quad+2E\bigg{[}((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))$
-$\quad\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\bigg{]}$
+$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
+$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
+$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(Y-\mu_{0}(D,X,1))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
+$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
+$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
+$\qquad+2E\bigg[((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))$
+$\qquad\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\bigg]$
 $=2E\left[((\mu_{0}(D,X,1)-\mu_{0}(D,X,0))((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))\}\right]$
-$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
-$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
+$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
+$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
 $=2E\left[((\mu_{0}(D,X,1)-\mu_{0}(D,X,0))((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))\}\right]$
-$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,1)-\mu_{0}(D,X,1))\mathbb{E}\left[\frac{Z}{p_{0}(D,X)}\Bigg{|}D,X\right]\right]$
-$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,0)-\mu_{0}(D,X,0))\mathbb{E}\left[\frac{(1-Z)}{1-p_{0}(D,X)}\Bigg{|}D,X\right]\right]$
-$\quad=0$
+$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,1)-\mu_{0}(D,X,1))\mathbb{E}\left[\frac{Z}{p_{0}(D,X)}\Bigg|D,X\right]\right]$
+$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,0)-\mu_{0}(D,X,0))\mathbb{E}\left[\frac{(1-Z)}{1-p_{0}(D,X)}\Bigg|D,X\right]\right]$
+$\qquad=0$
 and
-$\quad\partial_{r}\mathbb{E}[\tilde{\psi}_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
+$\qquad\partial_{r}\mathbb{E}[\tilde{\psi}_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big|_{r=0}$
 $=\mathbb{E}\left[\partial_{r}\tilde{\psi}_{2},(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big|_{r=0}\right]$
 $=\mathbb{E}\left[(\mu(D,X,1)-\mu_{0}(D,X,1))\right]-\mathbb{E}\left[(\mu(D,X,0)-\mu_{0}(D,X,0))\right]$
-$\quad-\mathbb{E}\left[\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
-$\quad+\mathbb{E}\left[\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
-$\quad-\mathbb{E}\left[\frac{Z(Y-\mu_{0}(1,D,X))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
-$\quad-\mathbb{E}\left[\frac{(1-Z)(Y-\mu_{0}(0,D,X))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]=0$
+$\qquad-\mathbb{E}\left[\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
+$\qquad+\mathbb{E}\left[\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
+$\qquad-\mathbb{E}\left[\frac{Z(Y-\mu_{0}(1,D,X))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
+$\qquad-\mathbb{E}\left[\frac{(1-Z)(Y-\mu_{0}(0,D,X))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]=0$
 since $\mathbb{E}[Z|D,X]=p_{0}(D,X)$, $\mathbb{E}[Z(Y-\mu_{0}(D,X,1))|D,X]=0$ and $\mathbb{E}[(1-Z)(Y-\mu_{0}(D,X,0))|D,X]=0$.
@@
 Equation (3.5) suggests the following score function for testing when $Z$ is multivalued discrete or continuous:
-$\quad\psi(W,\theta,\eta)$ (A.3)
+$\qquad\psi(W,\theta,\eta)$ (A.3)
 $=\sum_{l=1}^{L}(\mu(D,X,Z\in Z_{l})-\mu(D,X,Z\notin Z_{l}))^{2}$
@@
 $\mathbb{E}\Bigg[\sum_{l=1}^{L}(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))$
-$\quad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\Bigg{]}=0,$
+$\qquad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\Bigg]=0,$
 which holds by the same argument. Next, we show that Neyman orthogonality holds. First, note that
-$\quad\partial_{r}\mathbb{E}[\psi_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
+$\qquad\partial_{r}\mathbb{E}[\psi_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big|_{r=0}$
 $=\mathbb{E}\left[\partial_{r}\psi_{2},(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big|_{r=0}\right]$
 $=\sum_{l=1}^{L}\Bigg(\mathbb{E}\left[(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]-\mathbb{E}\left[(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
-$\quad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
-$\quad+\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
-$\quad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
-$\quad-\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]\Bigg{)}=0$
+$\qquad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
+$\qquad+\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
+$\qquad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
+$\qquad-\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]\Bigg)=0$
 since
-$\quad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
+$\qquad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
 $=\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]=0,$
 by the same arguments used before,
-$\quad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
+$\qquad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
 $=\sum_{l=1}^{L}\mathbb{E}\left[\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\Bigg|D,X\right]\right]$
@@
 $=\sum_{l=1}^{L}\mathbb{E}\left[(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right].$
-$\quad\partial_{r}\mathbb{E}[\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))]\big{|}_{r=0}$
+Further, we have
+$\qquad\partial_{r}\mathbb{E}[\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))]\big|_{r=0}$
 $=\mathbb{E}\left[\partial_{r}\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))\big|_{r=0}\right]$
-$=2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\right]$
-$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
-$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
-$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
-$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
-$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
-$\quad+2\sum_{l=1}^{L}\mathbb{E}\bigg{[}((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l})))$
-$\quad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))\cdot 1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))\cdot 1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\bigg{]}$
-$=2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\right]$
-$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]$
-$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
+$=2\sum_{l=1}^{L}\mathbb{E}\Big[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))$
+$\qquad\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\Big]$
+$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
+$\qquad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
+$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
+$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
+$\qquad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
+$\qquad+2\sum_{l=1}^{L}\mathbb{E}\bigg[((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l})))$
+$\qquad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))\cdot 1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))\cdot 1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\bigg]$
+$=2\sum_{l=1}^{L}\mathbb{E}\Big[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))$
+$\qquad\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\Big]$
+$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]$
+$\qquad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
 $=0.$
@@
 (2018) holds since
-$\quad\mathbb{E}[\psi(W,\theta_{0},\eta_{0})^{2}]$
+$\qquad\mathbb{E}[\psi(W,\theta_{0},\eta_{0})^{2}]$
 $\geq\mathbb{E}\left[\left(\sum_{l=1}^{L}\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)^{2}\right]$
@@
 by Assumption 6 and since $p_{l}(D,X)>c>0$ by construction for all $l=1,\dots,L$. This completes the proof.
+A.4 Figures
+Figure A1: Distribution of p-values: Oracle IV, Binary Simulations
+Figure A2: Distribution of p-values: Confounder $X_{1}$, Binary Simulations
+A.5 Tables
+Table A1: Application: All $\hat{S}$
+Panel A: Number of Prescriptions
+Candidate IV est se pval pct trimmed
+HH Income 0.003 0.038 0.929 0.386
+Random Assignment Instrument -0.007 0.010 0.472 0.001
+Living with a partner 0.013 0.009 0.163 0.335
+Ever Surveyed 0.013 0.009 0.143 0.294
+Any RX -0.019 0.011 0.078 0.391
+Need Dental Assistance Missing 0.060 0.032 0.063 0.905
+Application Received 0.024 0.013 0.061 0.381
+Employer Insurance 0.025 0.013 0.058 0.551
+Application Approved 0.040 0.015 0.007 0.681
+Surveyed All Months -0.540 0.176 0.002 0.771
+Retired 0.109 0.031 0.000 0.572
+Application Pending 0.057 0.016 0.000 0.703
+All Survey End 0.027 0.007 0.000 0.325
+OHP Insurance Missing 0.037 0.009 0.000 0.410
+Any Hospitalisation 0.057 0.014 0.000 0.656
+Panel B: Number of Doctor Visits
+Random Assignment Instrument -0.000 0.013 0.999 0.002
+Any Doctor Visit -0.002 0.011 0.871 0.344
+Application Approved 0.004 0.021 0.867 0.694
+Retired -0.027 0.073 0.705 0.601
+Application Received 0.010 0.015 0.496 0.398
+Need Dental Work Missing 0.039 0.052 0.447 0.917
+Ever Surveyed 0.011 0.011 0.287 0.315
+Surveyed in All Months -0.238 0.217 0.274 0.785
+Under 19 All Insured 0.026 0.016 0.111 0.475
+Ending Survey 0.023 0.012 0.064 0.343
+Application Pending 0.051 0.024 0.036 0.720
+HH Income Category 0.112 0.053 0.036 0.402
+Any Hospitalisation 0.064 0.025 0.012 0.669
+More than 30 hours employed 0.044 0.011 0.000 0.350
+Insurance Missing Variable 0.053 0.011 0.000 0.424
+OOP Prescription Costs -0.071 0.002 0.000 0.436
+OOP Medical Costs -0.060 0.001 0.000 0.429